### v1

Abstract
Federated learning is a machine learning method that supports training models on decentralized devices or servers, where each holds its local data, removing the need for data exchange. This approach is especially useful in healthcare, as it enables training on sensitive data without needing to share them. The nature of federated learning necessitates robust security precautions due to data leakage concerns during communication. To address this issue, we propose a new approach that employs selective encryption, homomorphic encryption, differential privacy, and bit-wise scrambling to minimize data leakage while achieving good execution performance. Our technique , FAS (fast and secure federated learning) is used to train deep learning models on medical imaging data. We implemented our technique using the Flower framework and compared with a state-of-the-art federated learning approach that also uses selective homomorphic encryption. Our experiments were run in a cluster of eleven physical machines to create a real-world federated learning scenario on different datasets. We observed that our approach is up to 90% faster than applying fully homomorphic encryption on the model weights. In addition, we can avoid the pretraining step that is required by our competitor and can save up to 20% in terms of total execution time. While our approach was faster, it obtained similar security results as the competitor.
Keywords:
Federated learning medical image datasets secure privacy-preserving machine learning
1 Introduction
Federated learning, a technique introduced by Google in 2016, offers a way to train machine learning models across decentralized data from mobile devices (federated, ). This approach has since gained widespread interest in both academic research and industry applications. By allowing model training on data that remains locally stored, federated learning helps protect privacy, making it an attractive solution for sensitive data environments. For example, in healthcare, it enables institutions to build shared machine learning models without the need to transfer patient data, thereby maintaining compliance with data protection laws. Overall, federated learning facilitates secure, privacy-conscious model training across decentralized networks.
The utilization of federated learning in sensitive domains such as healthcare necessitates robust security precautions due to its reliance on internet connections and the inherent risks of unauthorized data access and potential leaks. To address these privacy concerns, various security techniques have been integrated into federated learning frameworks. This research focuses on the implementation and evaluation of cryptology-based security precautions in federated learning.
This paper introduces a detailed experimental evaluation of different security techniques applied to federated learning, focusing on their impact on performance, computational overhead, and privacy protection. Specifically, we examine the effectiveness of homomorphic encryption, differential privacy, and a custom-developed fast and secure homomorphic encryption (FAS) technique. These methods were tested using different image datasets in a federated learning scenario simulated in a cloud environment.
Our key contributions are as follows:
Implementation of Security Techniques: We implemented secure aggregation methods, including homomorphic encryption, differential privacy, and FAS, within a federated learning model.
Development of FAS: We propose a novel FAS technique combining selective encryption, differential noise addition, and bitwise scrambling to achieve high security with minimal computational overhead.
Performance Evaluation: We provide a comparative analysis of encryption techniques, highlighting their overhead, scalability, and resilience against model inversion attacks, using metrics like MSSIM and VIFP.
Practical Insights: We assess the trade-offs between privacy, performance, and scalability, providing actionable insights for deploying secure federated learning systems in real-world applications.
The results demonstrate that FAS achieves a balance between robust security and computational efficiency, outperforming traditional techniques in scenarios where real-time performance and scalability are critical. The findings also highlight the suitability of FAS for resource-constrained environments like healthcare, where both privacy and efficiency are paramount.
The rest of this paper is organized as follows: The Datasets section describes the different image datasets used in our study. In Background on Deep Learning Models and Security Techniques, we discuss the deep learning models evaluated in our experiments, including ResNet-50 (resnet, ), DenseNet121 (densenet, ), EfficientNetB0 (efficientnet, ), MobileNet V2 (mobilenet, ), and elaborate on the security techniques applied.
The Experimental Setup for Federated Learning details our use of CloudLab (cloudlab, ) as the testbed and the Flower framework (flower, ) for federated learning experiments. The Evaluation section presents encryption overhead and time costs across datasets and models, with a detailed time comparison of security techniques.
In Comparison of Security Techniques, we evaluate differential privacy (differential, ), homomorphic encryption (gentry2009, ), and FAS, and compare these with similar works. The Practical Implications section discusses efficiency considerations and the balance between security and performance, followed by a Visual Comparison of FAS Approaches. Finally, we conclude with a summary of our findings and future research directions.
2 Background and Related work
Privacy-preserving techniques in federated learning have garnered significant attention due to the increasing need to protect sensitive data during model training across distributed devices. Federated learning, introduced by McMahan et al. (mcmahan2017communication, ), enables the training of models on decentralized data sources, but it also introduces vulnerabilities that could potentially leak private information through model updates.
Homomorphic encryption (HE) has emerged as a robust method for maintaining data confidentiality in federated learning. It allows computations to be performed directly on encrypted data, ensuring that sensitive information remains protected throughout the training process. Gentry’s foundational work on fully homomorphic encryption (FHE) (gentry2009, ) made arbitrary computations on encrypted data feasible. However, the high computational and storage overhead of FHE has limited its practical application, particularly in resource-constrained environments (boneh2006, ; brakerski2014efficient, ).
2.1 Efficiency Enhancements in Homomorphic Encryption
Several advancements have sought to improve the efficiency of homomorphic encryption:
Fan and Vercauteren’s BGV Scheme (fan2012somewhat, ): This scheme optimizes the bootstrapping process, making HE more practical by efficiently handling both addition and multiplication on ciphertexts. This approach has been critical in making HE viable for more complex computations in federated learning environments (smart2014fully, ). Xu et al. xu2019hybridalpha proposed HybridAlpha, an efficient privacy-preserving federated learning system that optimizes both privacy and performance. In comparison to our FAS approach, HybridAlpha leverages different privacy mechanisms such as differential privacy and secure aggregation to ensure efficient federated learning while maintaining data confidentiality.
Chillotti et al.’s TFHE Scheme (chillotti2016faster, ): TFHE introduced a faster bootstrapping process, making it highly efficient for secure Boolean computations, thus expanding the potential use of FHE in time-sensitive environments (chillotti2018faster, ).
Ducas and Micciancio’s FHEW Scheme (ducas2015faster, ): FHEW further improved the speed of homomorphic operations by optimizing the bootstrapping process, making FHE more feasible for larger-scale applications (ducas2015faster, ).
Cheon et al.’s CKKS Scheme (cheon2017homomorphic, ): This scheme supports approximate arithmetic operations on encrypted data, making it particularly efficient for machine learning and statistical analysis tasks where exact precision is not always required (cheon2018bootstrapping, ).
Brakerski, Gentry, and Vaikuntanathan’s BGV+ Scheme (brakerski2014efficient, ): This work built on the BGV scheme, introducing further optimizations that reduce the computational overhead of FHE, particularly in noise management, which is crucial for maintaining the accuracy of federated learning models under encryption (brakerski2014efficient, ).
2.2 Differential Privacy (DP) in Federated Learning
Differential privacy (DP) introduces controlled noise into the data or model updates, protecting individual data points while allowing accurate aggregate analysis. The foundational work by Dwork et al. (dwork2006calibrating, ) laid the groundwork for DP, and its application in deep learning was further developed by Abadi et al. (abadi2016deep, ), who integrated DP into the stochastic gradient descent (SGD) process. Subsequent research has focused on balancing the trade-off between privacy guarantees and model accuracy, with methods such as the moment accountant providing more precise privacy loss tracking (mironov2017renyi, ).
2.3 Selective Encryption in Federated Learning
Selective encryption provides a middle ground by encrypting only a subset of the model parameters, thereby reducing the computational burden while maintaining a reasonable level of security. Wu et al. (wu2019towards, ) demonstrated that selective encryption could significantly reduce overhead without compromising overall security by carefully choosing which parameters to encrypt. Other studies have explored similar strategies, focusing on optimizing the trade-off between encryption and efficiency in federated learning systems (li2019privacy, ; song2020privacy, ).
In this paper, we extend these concepts by introducing a novel selective parameter encryption scheme that incorporates differential noise and bitwise scrambling. This approach not only reduces the overhead associated with homomorphic encryption but also enhances privacy protection for the unencrypted parameters. By leveraging the efficiency of selective encryption and the principles of differential privacy, our method provides a practical solution for secure and efficient federated learning in sensitive domains.
2.4 Bitwise Scrambling in Federated Learning: Enhancing Security with Low Overhead
Bitwise scrambling is an obfuscation technique that further protects partly encrypted data by rearranging bits according to a cryptographic key, making it difficult for adversaries to extract meaningful information without access to the key. While scrambling techniques have long been used in secure communication protocols, recent research has adapted bitwise scrambling specifically for federated learning environments to enhance privacy with minimal computational overhead.
Yang et al. (yang2017bitwise, ) explored the use of bitwise scrambling as an additional layer of security in data transmission. By applying scrambling to unencrypted or selectively encrypted data, this technique provides a cost-effective way to obfuscate sensitive information during federated model training and aggregation. In federated learning, bitwise scrambling effectively prevents adversaries from performing direct analysis or reconstructing original data from partly encrypted models, even if the differential privacy noise is circumvented.
Other studies have further demonstrated the value of combining bitwise scrambling with homomorphic encryption in federated learning frameworks. For example, Halevi et al. (halevi2019faster, ) integrated scrambling with HE to enable secure, efficient transmission in distributed computing, showing that scrambling can effectively prevent chosen-ciphertext attacks by masking encrypted data patterns. This approach not only reduces the attack surface but also allows for faster processing times, as only a portion of the data requires full encryption while scrambled data remains obfuscated.
Bitwise scrambling, when combined with selective encryption and differential privacy as presented in this paper, enables a highly efficient privacy-preserving solution for federated learning. It obfuscates sensitive information without introducing significant additional computation, making it suitable for resource-constrained environments that still require robust security protections.
3 Motivation and Threat Model
3.1 Motivation
The growing adoption of federated learning (FL) in sensitive domains, such as healthcare and finance, has heightened the need for robust privacy-preserving mechanisms. While federated learning inherently minimizes the risk of data exposure by keeping raw data localized, it is still susceptible to a variety of privacy threats due to the exchange of model updates during training. Attackers can exploit these updates to infer sensitive information, such as patient health records or financial transactions, through techniques like model inversion or gradient leakage. These vulnerabilities necessitate enhanced security measures that preserve privacy without compromising performance.
Existing privacy-preserving techniques, such as full homomorphic encryption and differential privacy, have demonstrated effectiveness but face practical challenges. Fully homomorphic encryption ensures the highest level of data confidentiality but introduces significant computational and communication overhead, making it unsuitable for resource-constrained or time-sensitive environments. Differential privacy, while lightweight, can lead to degradation in model accuracy due to the introduction of noise, particularly in scenarios requiring high precision.
To address these challenges, this work introduces a novel FAS approach that balances security and efficiency. By selectively encrypting critical portions of the model updates and applying differential noise and bitwise scrambling to the remaining data, FAS reduces computational overhead while maintaining strong privacy guarantees. This method is especially valuable in federated learning scenarios where computational resources and latency are critical, such as medical imaging and real-time analytics.
The proposed approach provides a scalable, efficient, and robust solution to protect sensitive data in FL, bridging the gap between the security of full homomorphic encryption and the efficiency of lightweight techniques like differential privacy.
3.2 Threat Model
Federated learning (FL) systems are vulnerable to privacy threats due to the decentralized nature of training and the exchange of model updates. Adversaries can exploit these updates to reconstruct sensitive training data, such as medical images, or infer private attributes through techniques like gradient inversion and statistical analysis. This threat model assumes the presence of a semi-honest adversary who follows the protocol but attempts to infer sensitive information from intercepted or compromised data.
The adversary is capable of accessing model updates exchanged between clients and the central server and may compromise either the server or individual clients to gain insights. Potential attack vectors include gradient inversion attacks to reconstruct training data, membership inference to determine if specific data points were included in the training, and statistical analysis to extract private attributes. While all communications are assumed to be encrypted (e.g., using SSL), this does not prevent adversaries from analyzing the transmitted updates.
To mitigate these threats, our approach combines selective encryption, differential noise addition, and bitwise scrambling. Critical model parameters are encrypted, while unencrypted data is obfuscated with noise and scrambling to prevent reconstruction. Additionally, secure aggregation ensures that data remains protected throughout the learning process, providing robust defenses against privacy attacks without incurring excessive computational overhead. This layered strategy strikes a balance between security and efficiency, making it well-suited for sensitive applications like healthcare.
4 Proposed Model
This research introduces a novel FAS (fast and secure federated learning) approach tailored for secure and efficient federated learning. The model aims to strike a balance between robust privacy preservation and computational efficiency, addressing the limitations of fully homomorphic encryption (high overhead) and differential privacy (accuracy degradation). Our approach integrates three key components: selective encryption, differential noise addition, and bitwise scrambling, providing a multi-layered defense against privacy attacks while maintaining high performance.
4.1 Selective Encryption
Rather than encrypting the entire model, FAS selectively encrypts a fixed percentage of critical model parameters using a homomorphic encryption scheme. This subset is chosen uniformly, ensuring that sensitive components are protected without the computational burden of encrypting the entire dataset. By encrypting only a fraction of the data, FAS significantly reduces the overhead associated with full encryption while maintaining strong security for the most sensitive parameters.
4.2 Differential Noise Addition
For the remaining unencrypted parameters, differential noise is added to obfuscate their true values. The noise is carefully calibrated using a Laplacian mechanism, ensuring that individual data points cannot be inferred while preserving the overall utility of the model. This approach leverages the principles of differential privacy to provide an additional layer of protection for unencrypted data.
4.3 Bitwise Scrambling
To further enhance privacy, unencrypted parameters are subjected to bitwise scrambling using a cryptographic key. This process disrupts patterns in the data, making it computationally infeasible for adversaries to reconstruct the original values without access to the scrambling key. By combining scrambling with noise addition, the model ensures that unencrypted data is protected against both statistical analysis and direct reconstruction attacks.
4.4 Federated Learning Integration
The FAS model integrates seamlessly into the federated learning workflow. Each client encrypts and obfuscates its local model updates using the proposed techniques before transmitting them to the central server. The server aggregates the updates without requiring decryption, preserving privacy throughout the computation. The aggregated model is then securely shared with clients for the next training iteration.
4.5 Security and Efficiency Trade-Off
The proposed model achieves a balance between security and efficiency:
  • Security: By encrypting critical parameters and obfuscating the rest, FAS provides strong resistance to model inversion and membership inference attacks, with security results comparable to full homomorphic encryption.
  • Efficiency: The selective encryption strategy reduces computational and communication overhead by up to 90%, making it suitable for resource-constrained federated learning environments.
This layered approach ensures that sensitive data remains protected while enabling efficient model training, making FAS an effective solution for privacy-preserving federated learning in applications like healthcare and finance.
4.6 Privacy-Preserving Techniques
4.6.1 Homomorphic Encryption
Homomorphic encryption (HE) is a cryptographic technique that allows computations to be performed directly on encrypted data without decrypting it. This ensures that sensitive information remains confidential throughout the computation process.
In the context of federated learning, HE enables secure aggregation of model updates from distributed clients. The mathematical foundation of HE ensures that:
$\text{Dec}(\text{Eval}(\text{Enc}(x),\text{Enc}(y)))=f(x,y),$
where Enc and Dec denote encryption and decryption, respectively, and Eval represents a computation on encrypted data. This property is critical for privacy-preserving applications, as it prevents adversaries from accessing plaintext data (gentry2009, ).
4.6.2 Differential Privacy
Differential privacy (DP) is a privacy-preserving framework that introduces controlled noise to statistical computations, ensuring that the presence or absence of a single data point does not significantly influence the output.
Formally, a mechanism $M$ satisfies $\epsilon$-differential privacy if for any two adjacent datasets $D_{1}$ and $D_{2}$ differing in a single entry, and for any subset of outputs $O$:
$\frac{\Pr[M(D_{1})\in O]}{\Pr[M(D_{2})\in O]}\leq e^{\epsilon},$
where $\epsilon$ is the privacy budget, controlling the trade-off between privacy and utility (dwork2006calibrating, ).
In federated learning, DP is used to add noise to model updates or gradients before aggregation, protecting individual contributions while preserving the overall utility of the model. This makes DP a valuable tool for maintaining privacy in sensitive domains such as healthcare and finance.
4.7 Process Flow of the Proposed Model
The proposed model operates in three main stages: client-side processing, server-side processing, and client post-processing. Each stage has a set of steps, illustrated in pseudocode and explained in detail below.
4.8 Federated Learning System Process Flow
Figure 2 illustrates the client-side process, while Figure 3 shows the server-side process in the federated learning system with encryption and scrambling techniques.
Figure 1: General depiction of the federated learning model showing the processes of encryption, scrambling, and aggregation.
Figure 1 provides a general depiction of our federated learning model, demonstrating the encryption, scrambling, and noise addition steps across client nodes, and the subsequent aggregation performed by the server.
Figure 2: Client-side process in federated learning system: encryption, scrambling, and noise addition.
Figure 3: Server-side process in federated learning system: aggregation and re-scrambling.
The process begins at the client side, where local data is used to train a model, producing a weight vector. This weight vector is split, encrypted, and further obfuscated through bitwise scrambling and the addition of noise.
Once encrypted, the weight vectors are formatted to hexadecimal and transmitted to the server. The server then receives encrypted data from all clients and performs federated averaging (FedAvg) to aggregate the model parameters. Importantly, the aggregation is done on both encrypted and non-encrypted vectors, ensuring privacy is maintained while also allowing for more efficient computation.
After aggregation, the server mixes the results with a scrambling key before sending the updated parameters back to the clients, where they are decrypted and applied to the local models for further training.
This secure and privacy-preserving process enables effective model training across decentralized data while preventing unauthorized access to sensitive information.
4.8.1 Client-Side Processing
In the client-side processing stage, model weights ($w$) are prepared for secure transmission to the server by encrypting selected weights, obfuscating the remaining weights, and combining them into a single data structure. Initially, the weights are split into two subsets based on a predefined encryption percentage ($enc\_pct$) (Lines 6–8). The selected portion of the weights is encrypted using a homomorphic encryption function ($enc\_func$), and these encrypted weights are stored in $enc\_w$ (Lines 9–11). For the remaining weights, they are first formatted to resemble encrypted data ($fmt\_weight$), noise is added using $noise\_func$, and the noisy weights are scrambled with a cryptographic key ($scr\_key$) via a scrambling function ($scr\_func$) (Lines 12–16). These scrambled weights are stored in $scr\_w$. Finally, the encrypted and scrambled weights are combined into a single structure ($enc\_scr\_w$) (Line 17) and transmitted to the server (Line 18).
Algorithm 1 Client-Side Enc, Scrambling, and Noise Addition
1:
2:$w$: Model weights
3:$enc\_pct$: Percentage of weights to encrypt
4:$enc\_func$: Encryption function
5:$scr\_key$: Scrambling key
6:$scr\_func$: Scrambling function
7:$noise\_func$: Noise addition function
8:
9:$enc\_scr\_w$: Processed (encrypted and scrambled) weights
10:procedure Client Enc-Scr-Noise
11:     Split $w$ into encrypted and non-encrypted parts
12:     $n\leftarrow$ length of $w$
13:     $num\_enc\leftarrow(enc\_pct/100)*n$
14:     $enc\_w\leftarrow[]$
15:     $scr\_w\leftarrow[]$
16:     for $i=0$ to $num\_enc-1$ do
17:         $enc\_weight\leftarrow enc\_func(w[i])$
18:         Append $enc\_weight$ to $enc\_w$
19:     end for
20:     for $i=num\_enc$ to $n-1$ do
21:         $fmt\_weight\leftarrow$ format_as_encrypted($w[i]$)
22:         $noisy\_weight\leftarrow noise\_func($fmt_weight))))
23:         $scr\_weight\leftarrow scr\_func($noisy_weight,scr_key))))
24:         Append $scr\_weight$ to $scr\_w$
25:     end for
26:     Combine $enc\_w$ and $scr\_w$ into $enc\_scr\_w$
27:     Send $enc\_scr\_w$ to the server
28:end procedure
4.8.2 Server-Side Processing
On the server side, the received combined weights ($enc\_scr\_w$) are processed to aggregate updates securely without decryption. The scrambled weights are first unscrambled using the same cryptographic key ($scr\_key$) applied during client-side processing (Lines 5–7). Once unscrambled, these weights are aggregated using a processing function ($srv\_proc\_func$) (Lines 8–9). Homomorphic operations are applied directly on the encrypted weights without decryption (Line 9). After processing, the unscrambled weights are reformatted and scrambled again using $scr\_func$ and $scr\_key$ to preserve security (Lines 10–13). These processed and re-scrambled weights are then combined with the processed encrypted weights and sent back to the client (Line 14).
Algorithm 2 Server-Side Processing without Decryption
1:
2:$enc\_scr\_w$: Received weights from client
3:$scr\_key$: Scrambling key for re-scrambling
4:$srv\_proc\_func$: Server processing function
5:
6:$sec\_w$: Processed weights sent back to client
7:procedure Server Proc w/o Decrypt
8:     Receive $enc\_scr\_w$ from the client
9:     $proc\_w\leftarrow[]$
10:     $unscr\_w\leftarrow[]$
11:     for $w$ in $scr\_w$ do
12:         $unscr\_weight\leftarrow unscramble\_func(w,scr\_key)$
13:         Append $unscr\_weight$ to $unscr\_w$
14:     end for
15:     Process $unscr\_w$ using $srv\_proc\_func$
16:     Process $enc\_w$ using homomorphic operations (no decryption needed)
17:     Format $unscr\_w$ back to scrambled format:
18:     for $w$ in $unscr\_w$ do
19:         $re\_scr\_w\leftarrow scr\_func(w,scr\_key)$
20:         Append $re\_scr\_w$ to $proc\_w$
21:     end for
22:     Combine $proc\_enc\_w$ with $re\_scr\_w$
23:     Send $sec\_w$ back to the client
24:end procedure
4.8.3 Client Post-Processing
In the final stage, the client receives the processed weights ($sec\_w$) from the server (Line 2) and restores them for updating the local model. The encrypted subset of weights is decrypted using the appropriate decryption function (Lines 3–4), while the scrambled weights are unscrambled with the cryptographic key ($scr\_key$) to restore their original form (Lines 5–7). These two subsets of weights are then combined to reconstruct the full model weights (Line 8), which are subsequently used to update the client’s local model for the next round of federated learning (Line 9).
Algorithm 3 Client Post-Processing with Noise Restoration
1:
2:$sec\_w$: Weights received from server
3:$scr\_key$: Scrambling key for restoration
4:
5:$final\_w$: Updated model weights for next round
6:procedure Client Post-Proc with Noise Rest
7:     Receive $sec\_w$ from the server
8:     $proc\_enc\_w\leftarrow sec\_w[0:num\_enc]$
9:     $proc\_scr\_w\leftarrow sec\_w[num\_enc:]$
10:     for $w$ in $proc\_scr\_w$ do
11:         $restored\_w\leftarrow reverse\_scr\_func(w,scr\_key)$
12:         Append $restored\_w$ to $rest\_w$
13:     end for
14:     Combine $proc\_enc\_w$ and $rest\_w$ to form $final\_w$
15:     Proceed to next round of federated learning with $final\_w$
16:end procedure
4.9 Advantages of the Proposed Model
  • Enhanced Security: The combination of encryption, noise, and scrambling protects both encrypted and unencrypted data.
  • Efficiency: Reduces computational overhead by selectively encrypting only a portion of the weights.
  • Scalability: Optimized for large-scale federated learning systems with multiple clients.
4.10 Privacy Analysis
In this section, we formally show that our federated learning mechanism, which combines selective homomorphic encryption, differential privacy, and bitwise scrambling, maintains a well-defined differential privacy guarantee.
4.10.1 Setup and Definitions
We consider a federated learning setting with $n$ clients, each holding private data. The global model is described by parameters indexed by $[N]$. Let $S\subseteq[N]$ be the subset of parameters to be protected by selective homomorphic encryption (HE), and let $[N]\setminus S$ be the remaining parameters to which we add differential privacy (DP) noise.
Differential Privacy.
A mechanism $M$ is $\epsilon$-DP if for any two adjacent datasets $D,D^{\prime}$ (differing in one record) and any event $O$:
$\frac{\Pr[M(D)\in O]}{\Pr[M(D^{\prime})\in O]}\leq e^{\epsilon}.$
Selective Homomorphic Encryption.
Parameters in $S$ are encrypted using a semantically secure HE scheme. Semantic security ensures no polynomial-time adversary can distinguish ciphertexts of different messages, implying no additional privacy cost in a DP sense (0-DP).
Differential Privacy on the Remaining Parameters.
For each $i\in[N]\setminus S$, we add noise calibrated to $\epsilon_{i}$-DP. By the composition property of DP, releasing all these parameters together is $\left(\sum_{i\in[N]\setminus S}\epsilon_{i}\right)$-DP.
Bitwise Scrambling.
We define a scrambling function $\mathcal{T}_{k}:\mathcal{X}\to\mathcal{X}$, where $\mathcal{X}$ is the space of possible model outputs. $\mathcal{T}_{k}$ is deterministic, keyed by $k$, and independent of $D$ except through $M(D)$. This scrambling is a form of post-processing.
4.10.2 Privacy Guarantee
Theorem 4.1
Consider a mechanism $\mathcal{M}$ that on input dataset $D$:
  1. 1.
    Encrypts $W_{i}$ for $i\in S$ using a semantically secure HE scheme (gentry2009, ; Fontaine2007, ).
  2. 2.
    Adds noise to each $W_{j}$ for $j\in[N]\setminus S$ to achieve $\epsilon_{j}$-DP. Releasing all noisy parameters together is $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP (Dwork2014, ).
  3. 3.
    Applies a scrambling function $\mathcal{T}_{k}$ to the entire output, resulting in $\mathcal{M}^{\prime}(D)=\mathcal{T}_{k}(\mathcal{M}(D))$ (Dwork2006, ).
Then, $\mathcal{M}^{\prime}(D)$ is also $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP.
Proof
Step 1 (DP on $[N]\setminus S$): For $[N]\setminus S$, each parameter $W_{j}+\text{noise}_{j}$ is $\epsilon_{j}$-DP. By composition, releasing all these parameters is $\sum_{j\in[N]\setminus S}\epsilon_{j}$-DP (Dwork2014, ). Formally, for any adjacent $D,D^{\prime}$ and event $O$:
$\frac{\Pr[\mathcal{M}_{[N]\setminus S}(D)\in O]}{\Pr[\mathcal{M}_{[N]\setminus S}(D^{\prime})\in O]}\leq e^{\sum_{j\in[N]\setminus S}\epsilon_{j}}.$
Step 2 (Selective Encryption on $S$): Parameters in $S$ are encrypted with a semantically secure HE scheme. Without the secret key, no information about the plaintext is revealed, contributing 0-DP cost (gentry2009, ; Fontaine2007, ).
Step 3 (Scrambling as Post-Processing): The scrambling $\mathcal{T}_{k}$ is a deterministic, data-independent post-processing step. A fundamental property of DP is closure under post-processing: if $\mathcal{M}$ is $\epsilon$-DP, then $f(\mathcal{M})$ is also $\epsilon$-DP for any deterministic $f$ (Dwork2006, ; Dwork2014, ). Since $\mathcal{T}_{k}$ does not depend on $D$ beyond $\mathcal{M}(D)$ itself, applying $\mathcal{T}_{k}$ cannot increase or decrease the DP budget.
Combining these results:
$\frac{\Pr[\mathcal{M}^{\prime}(D)\in O]}{\Pr[\mathcal{M}^{\prime}(D^{\prime})\in O]}=\frac{\Pr[\mathcal{T}_{k}(\mathcal{M}(D))\in O]}{\Pr[\mathcal{T}_{k}(\mathcal{M}(D^{\prime}))\in O]}\leq e^{\sum_{j\in[N]\setminus S}\epsilon_{j}}.$
Thus, $\mathcal{M}^{\prime}(D)$ maintains the same DP guarantee as $\mathcal{M}(D)$.
5 Evaluation
In this section, we evaluate our federated learning setup using multiple security techniques applied to different image datasets. These techniques include homomorphic encryption, differential privacy, and our custom FAS. We assessed their effectiveness based on several metrics, such as test accuracy, communication cost, computational overhead, and the overall ability to maintain data privacy and security in a federated learning context.
The datasets were split into 10 equal-sized partitions and distributed among machines running Flower clients. Each machine trained its local model on its respective partition, with training occurring over 10 rounds of 20 epochs each. We used 10% of the training data to construct a validation set, ensuring consistent evaluation throughout the process. After the rounds were completed, the final model accuracy was measured on a separate test set for each dataset. Notably, all deep learning models were trained using only CPU resources to simulate realistic resource constraints.
The following tables compare the training times (in minutes) for different models across various datasets: CIFAR-10, CT Kidney, Lung, and diabetic retinophaty. Each table presents three encryption configurations:
5.1 Datasets used in experiments
To ensure a comprehensive evaluation of our model across various imaging contexts, we selected publicly available datasets with high usability ratings. Our criteria for dataset selection included broad coverage of diseases, utilization of diverse imaging techniques, and a substantial quantity of images for both training and validation purposes. With these requirements in mind, we chose the CIFAR-10 dataset for comparison, along with a medical dataset from Kaggle focusing on Lung Cancer Histopathological images. Below, we provide an overview of each dataset.
5.1.1 CIFAR-10 Dataset
The CIFAR-10 dataset is a widely recognized benchmark in image classification research. It contains 10 classes, each with 6,000 images of 32×32 pixels. The dataset includes 50,000 images for training and 10,000 images for testing, providing a diverse set of images for evaluating model performance. CIFAR-10 is frequently used for benchmarking the effectiveness of various models due to its relatively moderate complexity and diverse set of images (CIFAR10, ).
5.1.2 Chest X-ray (COVID-19, Pneumonia) Dataset
The Chest X-ray dataset includes 6,339 images classified into three categories: benign, COVID-19, and pneumonia, with 2,313 images per category. The grayscale X-ray images provide detailed views of internal body structures, assisting in the assessment of lung conditions. Representative images from each category are shown in Figure 4(ChestXray, )
Refer to caption Refer to caption Refer to caption
(a) Benign (b) COVID-19 (c) Pneumonia
Figure 4: Examples from the Chest X-ray (Covid, Pneumonia) dataset (mine, )
5.1.3 CT Kidney Dataset
The CT Kidney dataset consists of 12,446 CT scans in four categories: benign, cyst, stone, and tumor. The images are distributed as follows: 5,077 benign, 3,709 cyst, 1,377 stone, and 2,283 tumor cases. Each CT scan provides cross-sectional insights into the internal anatomy, aiding in diagnosis and analysis. Sample images are shown in Figure 5(kidney, )
Refer to caption Refer to caption
(a) Benign (b) Cyst
Refer to caption Refer to caption
(c) Stone (d) Tumor
Figure 5: Examples from the CT Kidney dataset (mine, )
5.1.4 Diabetic Retinopathy Detection Dataset
This dataset includes 88,645 fundus images, divided into two groups: benign and diabetic retinopathy. There are 65,342 benign and 23,303 diabetic retinopathy images. Fundus photography, utilized here, captures detailed views of the optic nerve, assisting in the detection of abnormalities. Sample images are displayed in Figure 6(diabetic, )
Refer to caption Refer to caption
(a) Benign (b) Diabetic retinopathy
Figure 6: Examples from the Diabetic Retinopathy Detection dataset (mine, )
5.1.5 Lung Cancer Histopathological Images
With 15,000 histopathological images across benign, squamous cell carcinoma (SCC), and adenocarcinoma (ACC) classes, this dataset provides 5,000 images per category. Sample images are shown in Figure 7.The data facilitates analysis and classification of histopathological images and was also cited in (mine, ).  (lungcolon, )
Refer to caption Refer to caption Refer to caption
(a) Benign (b) ACC (c) SCC
Figure 7: Examples from the Lung Cancer Histopathological Images dataset. (mine, )
In our study, we utilized multiple datasets to thoroughly assess model stability and performance under diverse conditions. The primary focus was on the CT Kidney Dataset, chosen for its consistency in delivering stable, reliable results. Other datasets, such as the Chest X-ray and Diabetic Retinopathy Detection datasets, provided complementary perspectives, allowing us to evaluate model generalizability across different medical imaging types. The CT Kidney dataset’s robustness was especially beneficial, enabling us to prioritize security technique evaluation over model-specific tuning.
6 Experimental Setup for Federated Learning
In this section, we present the experimental setup used to evaluate the performance and security of various cryptographic techniques in a federated learning context on medical image datasets.
6.1 CloudLab Environment
All experiments were conducted on CloudLab (cloudlab, ), a cloud computing testbed designed for systems research. Our setup utilized eleven physical machines in a shared-nothing cluster, connected via Gigabit Ethernet (10 Gbps) to form a local area network. Each machine was equipped with two Intel E5-2683 v3 (14-core) CPUs at 2.00 GHz, 256 GB RAM, and two 1 TB hard drives. Each core had two hardware threads, providing ample processing power for federated learning tasks.
6.2 Federated Learning Framework: Flower
Flower (flower, ) is an open-source federated learning framework, compatible with popular machine learning libraries such as PyTorch (pytorch, ), TensorFlow (tensorflow, ), and MXNet (mxnet, ). In our experiments, we used TensorFlow to train the models and assess accuracy. The Flower framework provided flexible APIs and allowed us to implement various federated optimization strategies. One of the CloudLab machines was designated as the Flower server, while the remaining ten served as Flower clients.
6.3 Federated Learning Setup
Each dataset was divided into 10 equal-sized partitions and distributed across the Flower clients. The centralized federated learning setup consisted of 10 clients and a single server, simulating a typical federated learning framework. Each model was trained over 10 rounds, with each round consisting of 20 epochs. A validation set comprising 10% of the training data was used to monitor model performance. The final model accuracy was evaluated on the test set after completing all rounds.
This controlled environment allowed us to systematically examine the performance and security of different cryptographic methods on decentralized data. Only CPUs were utilized in the setup, creating a realistic resource-constrained environment for federated learning.
6.4 Rationale for Excluding Decentralized Federated Learning
In this study, we opted for a centralized federated learning setup instead of a decentralized approach. A primary consideration was the lack of mutual trust among clients. In decentralized federated learning, clients typically exchange data or model updates directly with one another, creating a network without a central coordinator. However, in our case, clients are unwilling to share their data directly with other participants due to confidentiality concerns and the sensitivity of the information involved.
Centralized federated learning provides a more controlled environment where a central server coordinates the aggregation of model updates, allowing clients to keep their data strictly local while still benefiting from collaborative training. This centralized coordination ensures that data privacy is maintained between clients, as each client only communicates with the trusted server, avoiding direct data exchange with other clients. This setup aligns better with the privacy requirements of our application, where maintaining strict data confidentiality is paramount.
6.5 Deep Learning Models
Deep learning has revolutionized disease diagnosis using medical image datasets, with various convolutional neural networks (CNNs) demonstrating success since 2014. In this study, we utilized five popular models—ResNet-50, DenseNet121, EfficientNetB0, MobileNet V2 for broader evaluation.
6.5.1 ResNet-50
ResNet-50 mitigates the vanishing gradient problem using skip connections, enabling effective training of deep networks and achieving high accuracy in complex tasks.
6.5.2 DenseNet121
DenseNet121 connects each layer to all previous layers, enhancing feature reuse and propagation for efficient, accurate performance in resource-constrained setups.
6.5.3 EfficientNetB0
EfficientNetB0 optimizes accuracy and efficiency by balancing network width, depth, and resolution, making it ideal for computationally constrained scenarios.
6.5.4 MobileNet V2
MobileNet V2 reduces computational demands with lightweight convolutions, maintaining high accuracy for applications on mobile and embedded devices.
- Non Enc: The model is trained without any encryption.
- Full Enc: All data and model weights are fully encrypted using homomorphic encryption.
- Partly Enc: A partly encryption approach is used, encrypting only a portion of the data while scrambling the rest for security.
The overhead columns represent the additional time required for encryption relative to the non-encrypted baseline.
Full Overhead captures the additional time incurred by full encryption, while
Partly Overhead shows the time overhead introduced by partly encryption.
Table 1: Model Parameters and Data Sizes (Fully Encrypted (Full Enc) and Partly Encrypted (Partly Enc))
Model Params (M) Unenc Size (MB) Full Enc (GB) Partly Enc (GB)
EffNetB0 5.3 21.2 2.71 0.281
DenseNet121 7.0 28.0 3.58 0.487
MobileNetV2 3.4 13.6 1.79 0.156
ResNet-50 25.6 102.4 12.4 1.64
Table 1 presents a comparison of model parameters and data sizes for both fully and partly encrypted models. The ”Params” column lists the number of model parameters in millions, while the ”Unenc Size” and ”Enc Size” columns show the data size for unencrypted, fully encrypted, and partly encrypted models. Partly encryption significantly reduces the size of encrypted data compared to full encryption, as illustrated by EfficientNetB0, where the fully encrypted size is 2.71GB, and the partly encrypted size is only 0.281GB.
Table 2: CIFAR-10 Performance Comparison (Time in Minutes)
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 8m 111m 21m 103m 13m
DenseNet 11m 219m 35m 208m 24m
MobileNet 7m 68m 15m 61m 8m
ResNet-50 16m 530m 68m 514m 52m
Table 2 highlights training times for different models on the CIFAR-10 dataset. Fully encrypted models exhibit significantly longer training times than partly encrypted models due to the overhead of encrypting the entire dataset. For example, DenseNet121 requires 219 minutes to train when fully encrypted, compared to only 35 minutes with partly encryption, illustrating the time-saving benefits of selective encryption without sacrificing data security.
Table 3: Performance on CT Kidney Dataset
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 112m 228m 127m 116m 15m
DenseNet121 301m 514m 326m 213m 25m
MobileNetV2 91m 163m 100m 72m 9m
ResNet-50 351m 861m 399m 510m 48m
Table 3 shows training times for models on the CT Kidney dataset. The overheads are consistent with the number of model parameters, with heavier models incurring more time overhead. The training overheads for full and partly encryption across models are largely in line with what is expected based on the size and complexity of the models.
Table 4: Performance on Lung Dataset
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 241m 355m 254m 114m 13m
DenseNet121 287m 506m 312m 219m 24m
MobileNetV2 225m 467m 231m 242m 6m
ResNet-50 319m 834m 372m 515m 53m
Table 4 focuses on the Lung dataset, where training times for encrypted and non-encrypted models are compared. As seen with the CIFAR-10 and CT Kidney datasets, larger models such as ResNet-50 incur the most substantial encryption overhead, while smaller models like MobileNetV2 and EffNetB0 see lighter impacts. The results once again illustrate that encryption overhead correlates directly with the number of parameters and the model’s complexity.
Table 5: Performance on diabetic retinophaty Dataset
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 578m 698m 591m 120m 13m
DenseNet121 1298m 1542m 1351m 244m 53m
MobileNetV2 546m 610m 555m 64m 9m
ResNet-50 622m 1140m 673m 518m 51m
Table 5 displays the performance on the diabetic retinophaty dataset, which follows the same trend as the previous datasets. Models with more parameters, like ResNet-50, experience higher encryption overheads. The consistency of overhead across different datasets suggests that encryption overheads scale predictably with model size and parameter count, as expected.
6.6 Graphical Comparisons and Analysis
Figure 8: Encryption Overhead Comparison Between Different Models
Figure 8 compares the encryption overheads between fully encrypted and partly encrypted models, revealing a significant reduction in overhead with partly encryption. Partly encryption consistently reduces overhead by approximately 90%, providing a highly efficient method for securing models while still maintaining robust security. This balance between drastically reduced encryption time and maintaining data protection makes partly encryption a compelling alternative to full encryption, especially for larger models, where the time savings are particularly pronounced.
Figure 9: Encrypted file size per encryption level (10% - 100%)
Figure 9 illustrates the communication costs associated with different encryption levels for various models, including EfficientNetB0, DenseNet121, MobileNetV2, ResNet-50. The X-axis represents the percentage of encryption applied to the model weights, ranging from 10% to 100% (encryption). The Y-axis displays the corresponding data size in gigabytes (GB), reflecting the communication cost for each model.
The graph shows that as the encryption percentage increases, so does the communication cost. For instance, EfficientNetB0 starts at an unencrypted size of 21.2 MB and grows to 2.71 GB at 100% encryption. Partly encryption mitigates this increase, with significantly lower communication costs at lower encryption levels, such as 0.281 GB for partly encrypted EfficientNetB0.
Across all datasets, fully encrypted models exhibit higher communication costs and computational overhead. Partly encryption, however, reduces both metrics considerably, making it a more practical option for real-time applications where security and performance need to be balanced.
Figure 10: Accuracy Vs Encryption Percentage Across Different Models
Figure 10 demonstrates that model accuracy remains robust even as the encryption percentage decreases from 90% to 10%. The minimal fluctuations (less than 1%) in accuracy suggest that the encryption technique maintains performance while offering strong security. Any slight variations can be attributed to noise or minor differences in the training process. This experiment was conducted in an offline environment, rather than a federated setup, to isolate and clearly show the effects of encryption on model accuracy. dataset used in this accuracy representation is lung cancer dataset.
Summary: Accuracy remains largely unaffected by partly encryption, maintaining a stable performance across different models. The minimal decrease in accuracy demonstrates that partly encryption provides a viable alternative to full encryption, offering security without compromising model effectiveness.
Figure 11: Training Time and Encrypted Size Comparison for Fully and Partly Encrypted Models For CT Kidney dataset
Figure 11 compares the training times and encrypted sizes of both fully and partly encrypted models for CT Kidney dataset. The graph visually illustrates the significant reduction in encryption overhead when applying partly encryption techniques compared to full encryption.For larger models like ResNet-50, fully encrypted models show a sharp increase in both training time and encrypted size.
FAS delivers substantial improvements in performance efficiency, reducing training times and encrypted sizes across a variety of datasets. It emerges as a viable strategy for federated learning where computational resources are limited, yet security remains a top priority.
6.7 Effects on Security
Before conducting our main security experiments, we implemented several security tests to evaluate the robustness of our encryption methods. These tests included integrity checks, timing attack tests, noise and precision evaluations, and simulations of chosen-ciphertext and chosen-plaintext attacks. Additionally, differential attack tests and statistical distribution checks were performed to ensure that encryption did not compromise data integrity or privacy. Based on these initial tests, we found that 10% partly encryption successfully passed the same security tests as 100% full encryption, indicating that partly encryption maintains a high level of security while reducing computational demands.
Table 6: Full vs Partly Encryption: Security Test Results
Test Partly Enc. Full Enc.
Integrity Check False False
Chosen-Plaintext Attack True True
Chosen-Ciphertext Attack True True
Noise and Precision Test True True
Timing Attack Test True True
Differential Attack Test True True
Statistical Dist. Test True True
Homomorphism Test True True
Leakage Test True True
Table 6 presents the results of comprehensive security tests comparing partly encryption with scrambling to full encryption. Both encryption schemes underwent a variety of cryptographic attacks and precision tests to evaluate their robustness. The results indicate that partly encryption offers comparable security outcomes to full encryption across most criteria, without significant compromises in protection.
6.7.1 Evaluating Privacy with MSSIM and VIFP
Model inversion attacks threaten federated learning by reconstructing training data from model updates. To evaluate privacy robustness, we utilize the Mean Structural Similarity Index (MSSIM) and Visual Information Fidelity in the Pixel domain (VIFP).
MSSIM quantifies structural similarity between original and reconstructed images, with low scores indicating effective privacy measures. Similarly, VIFP assesses visual quality based on human perceptual models, detecting fine-grained artifacts and the impact of techniques like differential privacy noise and bitwise scrambling. Both metrics are widely recognized for their ability to evaluate privacy in adversarial scenarios.
Our results show that selective encryption combined with scrambling and noise significantly degrades reconstruction quality, with MSSIM stabilizing at 52% for 20% encryption and VIFP confirming similar trends across multiple encryption levels. These findings highlight the effectiveness of our approach and the utility of MSSIM and VIFP as reliable privacy evaluation metrics.
Following these tests, we further evaluated the effectiveness of partly encryption against model inversion attacks. In these attacks, we attempted to reconstruct images from the encrypted models and compared the reconstructed images to the originals using VIFP (vifp, ) and MSSIM (mssim, ) metrics.
Despite using only 10% encryption, the reconstructed images displayed significant distortion. For instance, at 10% encryption, the MSSIM score was 58%, indicating substantial deviation from the original image. As the encryption percentage increased, the score stabilized at around 52%, further demonstrating that partly encryption provides strong protection against inversion attacks, even with reduced encryption levels.
Figure 12: MSSIM Scores across Encryption Percentages for Various Models
Figure 12 illustrates the MSSIM scores for various models across different levels of partly encryption. The displayed scores were obtained by averaging results from multiple experimental runs to account for randomness in selective encryption, scrambling, and noise. As encryption increases from 10%, there is a sharp decline in MSSIM scores, with most models stabilizing around 20% encryption. These results suggest that even low levels of partly encryption can achieve security comparable to full encryption, with significantly reduced computational demands. Notably, at 100% encryption, the resistance against inversion attacks is very close to the resistance achieved at 10% and similar to that at 20% encryption, highlighting the effectiveness of selective encryption even at lower levels.
Figure 13: VIFP Scores across Encryption Percentages for Various Models
Figure 13 shows VIFP scores across different encryption percentages for various models. As with MSSIM, VIFP scores drop sharply at 10% encryption and stabilize at around 20%. The displayed scores were repeated across multiple experimental runs to ensure robustness and mitigate the impact of randomness introduced by scrambling and noise. Importantly, the resistance to model inversion attacks at 100% encryption** is very similar to that achieved at 10% and 20% encryption levels. This stability indicates that partly encryption offers effective protection against model inversion attacks, even at lower encryption levels.
6.8 Cross-testing with related work
6.8.1 Comparison with MASKCRYPT (10506637, )
MASKCRYPT is a selective encryption method designed to minimize the overhead of fully encrypting model updates by applying a gradient-guided encryption mask. Instead of pre-training, MASKCRYPT calculates gradient values during each round, determining which model weights are critical for privacy protection and selectively encrypting those. This reduces communication overhead by up to 4.15× compared to full model encryption and decreases membership inference attack accuracy to 49.2% (ma2020secure, ). However, MASKCRYPT’s gradient mask is not always optimized for selecting the most sensitive gradients, which can result in lower initial security scores.
Figure 14: MSSIM and VIFP scores over rounds : MASKCRYPT
Figure 14 depicts the VIFP scores for MASKCRYPT over five rounds of federated learning. The MASKCRYPT model was recreated by us to evaluate its performance. While the results may not fully match the original models values, they offer a reliable insight into MASKCRYPT’s behavior. As observed in the graph, MASKCRYPT requires multiple rounds to achieve an acceptable encryption stability level, with a noticeable reduction in VIFP scores between rounds 1 and 3, eventually stabilizing around round 5.
This gradual improvement contrasts with our proposed FAS method, which stabilizes at a much earlier stage. This difference in behavior further highlights the efficiency of our approach in achieving secure and stable encryption without requiring as many iterations.
In contrast, our FAS method provides a more efficient balance between security, performance, and scalability. By combining selective homomorphic encryption with differential noise and bitwise scrambling, our approach encrypts a fixed percentage of weights (e.g., 10%) without the need for gradient sensitivity analysis or pre-training. This allows us to achieve high computational efficiency, with a 90% reduction in processing time compared to full homomorphic encryption, while also providing robust resistance to model inversion attacks. The layered privacy provided by noise and scrambling further enhances security without incurring the extra communication and computational costs associated with MASKCRYPT’s per-round gradient analysis process. Importantly, because our method doesn’t require per-round gradient recalculations or client-specific setups, it scales effectively across larger federated networks without additional setup costs for each client.
6.8.2 Comparison with FedML-HE (jin2024fedmlheefficienthomomorphicencryptionbasedprivacypreserving, )
FedML-HE employs a targeted encryption strategy in federated learning by pre-training the model to identify critical weights based on gradient sensitivity, which are then selectively encrypted.FedML-HE model was recreated by us to evaluate its performance. While the results may not fully match the original models values, they offer a reliable insight into FedML-HE’s behavior. While this approach offers strong privacy protection for the most sensitive weights, it introduces notable time and resource costs due to the pre-training phase. Additionally, in FedML-HE, each client independently generates an encryption mask on its local dataset, which is then aggregated at the central server. While this avoids directly sharing a mask across clients, it can limit scalability, as each client’s mask may vary based on smaller, locally available datasets, which may lead to inconsistencies in aggregated model accuracy.
Our FAS method avoids these limitations by providing privacy without the need for pre-training or mask aggregation. Instead of selecting weights through gradient sensitivity, our method encrypts a fixed percentage of weights and augments this with differential noise and bitwise scrambling. This eliminates the computational burden of FedML-HE’s pre-training phase and simplifies scalability, as our approach does not depend on client-specific training data characteristics. At 10% encryption, our approach achieves comparable security results (e.g., MISSIM score of 58%) to FedML-HE’s pre-trained encryption (52%) but with significantly reduced complexity, better scalability, and faster performance.
In summary, our FAS method provides a more efficient and privacy-preserving alternative to FedML-HE, particularly suited for federated learning environments where computational resources and data confidentiality are key considerations. By reducing the computational burden and ensuring consistency across clients without needing mask aggregation, FAS is easily scalable to large federated networks.
Figure 15: Comparison of MSSIM Scores: Our work vs. FedML-HE vs. MASKCRYPT
Figure 16: Comparison of VIFP Scores: Our work vs. FedML-HEvs. MASKCRYPT
Figures 15 and 16 provide a visual comparison of our FAS method, FedML-HE, and MASKCRYPT. In Figure 15, we observe that at 10% encryption, FedML-HE demonstrates slightly higher stability with a 52% MSSIM score compared to our 58%. MASKCRYPT, on the other hand, achieves a 55% MSSIM score, placing it between the two. However, both FAS and MASKCRYPT converge with FedML-HE around 20% encryption, resulting in nearly identical security performance across all methods. Figure 16 similarly shows that while FedML-HE achieves marginally better stability at lower encryption levels, and MASKCRYPT performs slightly better than FAS in these cases, our FAS technique provides comparable security with significantly reduced computational costs. MASKCRYPT incurs a higher computational overhead due to its sensitivity mask recalculation in every round, making it less efficient in scenarios requiring real-time performance. This efficiency gap makes FAS an ideal choice for scalable and privacy-sensitive applications, particularly in real-time federated learning scenarios.
In conclusion, each method has unique advantages based on application needs. For cases where lower encryption levels suffice and computational time is less critical, FedML-HE or MASKCRYPT may offer a slight edge in stability. However, for environments that demand higher encryption percentages along with performance and scalability, our FAS approach provides equivalent security with reduced time costs. This makes FAS a more efficient and scalable choice for federated learning, particularly in applications that prioritize real-time performance.
Table 7: Comparison of Encryption Techniques for Different Models
Method Model Non Enc total (m) Overhead (m)
FAS EffNetB0 56 69 13
FEDML-HE EffNetB0 56 99 43
MASKCRYPT EffNetB0 56 89 33
FAS DenseNet121 152 176 24
FEDML-HE DenseNet121 152 210 58
MASKCRYPT DenseNet121 152 200 48
FAS MobileNetV2 46 52 6
FEDML-HE MobileNetV2 46 69 23
MASKCRYPT MobileNetV2 46 64 18
FAS ResNet-50 176 224 48
FEDML-HE ResNet-50 176 278 102
MASKCRYPT ResNet-50 176 246 70
Table 7 shows the non-encrypted, partly encrypted, and overhead times for different models using three encryption methods: Our FAS, FedML-HE, and MASKCRYPT. The data in the table is a logical reconstruction of the models based on the methodologies described in the respective papers, and may not exactly represent the original implementations. Our FAS consistently has the lowest overhead across models, making it the most efficient. FedML-HE shows the highest overhead due to pre-training, while MASKCRYPT offers a balance but initially has lower security, catching up after several rounds. Overall, our FAS is the best option for efficiency and scalability in federated learning.
6.9 Comparative Analysis of Encryption Techniques Across Datasets
This section provides a detailed comparison of the three encryption techniques—FAS, FEDML-HE, and MASKCRYPT—across five datasets: CIFAR-10, Diabetic Retinopathy, COVID, Lung, and Kidney datasets. The analysis emphasizes the efficiency of FAS in reducing both overhead and encryption time compared to the other methods.
6.9.1 CIFAR-10 Dataset
Figure 17 demonstrates the performance on the CIFAR-10 dataset. FAS consistently achieves the lowest total training times and overhead across all models. Unlike FEDML-HE, which incurs significant pretraining costs for sensitivity mask creation, and MASKCRYPT, which generates masks at each round, FAS employs a random masking strategy with minimal computational burden. This positions FAS as the fastest and most efficient method for CIFAR-10.
6.9.2 Diabetic Retinopathy Dataset
As depicted in Figure 18, FAS significantly reduces overhead times for the diabetic retinopathy dataset compared to FEDML-HE and MASKCRYPT. The pretraining in FEDML-HE and the per-round mask generation in MASKCRYPT add substantial time costs, making FAS the preferred choice for this dataset.
6.9.3 COVID Dataset
The results for the COVID dataset are presented in Figure 19. FAS demonstrates a clear advantage, particularly in lightweight models like MobileNetV2, where its simplicity translates to much lower overhead times. The benefits of avoiding sensitivity mask-related computations are most apparent in this dataset.
6.9.4 Lung Dataset
Figure 20 showcases the performance on the lung dataset. FAS continues to outperform the other techniques by a significant margin, particularly in larger models such as ResNet-50. This demonstrates its scalability and adaptability to more complex models and datasets.
6.9.5 Kidney Dataset
Finally, Figure 21 highlights the results for the kidney dataset. Across all models, FAS achieves the best performance, confirming its robustness and suitability for diverse datasets. The elimination of pretraining and per-round computations makes FAS a practical and efficient solution.
Figure 17: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (CIFAR-10).
Figure 18: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (Diabetic Retinopathy).
Figure 19: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (COVID).
Figure 20: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (Lung).
Figure 21: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (Kidney).
Figures 17, 19, 18, 21 and 20 shows that across all datasets, FAS consistently outperforms FEDML-HE and MASKCRYPT in terms of total training time and overhead. The absence of sensitivity mask pretraining or extra per-round computations is the key factor contributing to its superior performance. These results establish FAS as the optimal choice for applications requiring fast and efficient encryption without compromising security. We can see that in smaller datasets which training time is not significant, difference between encryption technique becomes more obvious.
6.10 Handling Data Skew in Encryption-Based Techniques
In this section, we explore the impact of data skew on different encryption techniques, focusing on the robustness of the proposed FAS method compared to existing approaches such as MASKCRYPT and Fedml-HE.
6.10.1 Challenges with Mask-Based Techniques
Both MASKCRYPT and Fedml-HE rely on sensitive masks to select important gradients for encryption. These masks are designed to prioritize gradients that have a significant impact on the model’s accuracy. However, in scenarios with skewed data distributions, the importance of gradients becomes difficult to determine accurately. As a result, the encryption decisions made by these methods often degrade to levels resembling random encryption. This reduces their effectiveness, as observed in the MSSIM and VIFP score comparisons.
6.10.2 Robustness of the FAS Technique
In contrast, the proposed FAS technique, which operates as a random encryption method, is unaffected by data skew. Unlike mask-based techniques, FAS does not depend on the model’s accuracy or gradient importance for its encryption process. Instead, it employs a combination of encryption, scrambling, and noise addition. These components ensure that the method maintains consistent performance irrespective of the underlying data distribution. The robustness of this approach makes it particularly suitable for scenarios where data skew is prevalent, such as federated learning environments with heterogeneous clients.
6.10.3 Experimental Results and Comparisons
Figure 22: MSSIM and VIFP score comparison for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 23: MSSIM and VIFP score comparison for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 24: MSSIM and VIFP score comparison on lung dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 25: MSSIM and VIFP score comparison on lung dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 26: MSSIM and VIFP score comparison on kidney dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 27: MSSIM and VIFP score comparison on kidney dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 28: MSSIM and VIFP score comparison on covid dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 29: MSSIM and VIFP score comparison on covid dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 30: MSSIM and VIFP score comparison on diabetic retinopathy dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 31: MSSIM and VIFP score comparison on diabetic retinopathy dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figures 22, 23, 24, 25, 26, 27, 28, 29, 30, 31 shows that the proposed FAS technique, with its reliance on encryption, scrambling, and noise, demonstrates superior robustness to data skew across all evaluated datasets—Kidney, Lung, COVID, and Diabetic Retinopathy—compared to existing gradient-based encryption methods. This property ensures that FAS retains its performance and security benefits even in diverse and challenging real-world scenarios. The results from experiments using the EfficientNetB0 model across these datasets highlight that the accuracy of FAS is minimally impacted by data skew, unlike other methods, which exhibit noticeable performance degradation under similar conditions. This consistent robustness across datasets underscores the adaptability and reliability of the FAS technique for privacy-preserving machine learning.
  • For MASKCRYPT and Fedml-HE, the MSSIM and VIFP scores under skewed data are significantl degraded, closely resembling random encryption.
  • In contrast, the FAS technique maintains reduced degregation across all conditions, as its random encryption approach is independent of the data distribution.
  • The control value has been kept same with skewed and normal to show the random selective encryption comparison.
Effect of Data Skew on Training Performance
Figure 32: Effect of Data Skew on CIFAR Dataset Across Different Techniques
Figure 32 illustrates the uniform reduction in accuracy across three techniques—FAS, MASKCRYPT, and Fedml-HE—when training on skewed data using the CIFAR dataset. While all techniques show similar accuracy drops under skewed conditions, the implications for their security mechanisms differ significantly.
  • MASKCRYPT and Fedml-HE: Both techniques rely on sensitivity masks to ensure data security. As training accuracy decreases under skewed conditions, the quality and reliability of these sensitivity masks are compromised, potentially leaving the system insecure until the mask stabilizes.
  • FAS: Unlike the other techniques, FAS does not depend on sensitivity masks. This independence ensures that the MS-SSIM and VIFp scores of FAS remain stable, even when training on skewed data. The inherent robustness of FAS allows it to maintain its security guarantees regardless of data distribution.
These results highlight a critical advantage of FAS: its ability to decouple data security from training accuracy. While longer training epochs can partially recover accuracy for MASKCRYPT and Fedml-HE, their reliance on sensitivity masks introduces a window of vulnerability during the calibration phase. In contrast, FAS maintains consistent security and performance, making it a more robust choice under challenging data conditions.
7 Conclusion
This paper evaluated privacy-preserving techniques in federated learning, focusing on different datasets and comparing differential privacy, homomorphic encryption, and our custom FAS approach.
Full homomorphic encryption offers the highest data protection but incurs significant computational costs, making it suitable only for scenarios prioritizing confidentiality over performance. Differential privacy provides lightweight privacy with minimal impact on computation, ideal for moderate security requirements. Our FAS method strikes a balance by achieving security comparable to full encryption while significantly reducing training time and overhead, making it efficient for large-scale federated learning and resource-constrained environments.
FAS’s layered approach—combining selective encryption, bitwise scrambling, and differential noise—demonstrates strong resilience against model inversion attacks without requiring pre-training or complex mask aggregation, outperforming FedML-HE and MASKCRYPT in scalability and efficiency. FAS offers slighly better security compared to models that require accurate sensitivity masks for data skews or operate under general low-accuracy conditions.
In summary, FAS offers an effective middle ground, balancing security and performance for real-time, privacy-sensitive applications like healthcare. Future research will refine these techniques and explore hybrid approaches across diverse datasets and federated environments to enhance scalability and applicability.
8 Acknowledgments
The first author (A. K.) was supported by the Republic of Türkiye.