### v2

# A Selective Homomorphic Encryption Approach for Faster Privacy-Preserving Federated Learning

Abstract
Federated learning is a machine learning method that supports training models on decentralized devices or servers, where each holds its local data, removing the need for data exchange. This approach is especially useful in healthcare, as it enables training on sensitive data without needing to share them. The nature of federated learning necessitates robust security precautions due to data leakage concerns during communication. To address this issue, we propose a new approach that employs selective encryption, homomorphic encryption, differential privacy, and bit-wise scrambling to minimize data leakage while achieving good execution performance. Our technique FAS (fast and secure federated learning) is used to train deep learning models on medical imaging data. We implemented our technique using the Flower framework and compared with a state-of-the-art federated learning approach that also uses selective homomorphic encryption. Our experiments were run in a cluster of eleven physical machines to create a real-world federated learning scenario on different datasets. We observed that our approach is up to 90% faster than applying fully homomorphic encryption on the model weights. In addition, we can avoid the pretraining step that is required by our competitor and can save up to 46% in terms of total execution time. While our approach was faster, it obtained similar security results as the competitor.
Index Terms:
Federated learning medical image datasets secure privacy-preserving machine learning
I Introduction
Federated learning, a technique introduced by Google in 2016, offers a way to train machine learning models across decentralized data from mobile devices [1]. This approach has since gained widespread interest in both academic research and industry applications. By allowing model training on data that remains locally stored, federated learning helps protect privacy, making it an attractive solution for sensitive data environments. For example, in healthcare, it enables institutions to build shared machine learning models without the need to transfer patient data, thereby maintaining compliance with data protection laws. Overall, federated learning facilitates secure, privacy-conscious model training across decentralized networks.
The utilization of federated learning in sensitive domains such as healthcare necessitates robust security precautions due to its reliance on internet connections and the inherent risks of unauthorized data access and potential leaks. To address these privacy concerns, various security techniques have been integrated into federated learning frameworks. This research focuses on the implementation and evaluation of cryptology-based security precautions in federated learning.
This paper introduces a detailed experimental evaluation of different security techniques applied to federated learning, focusing on their impact on performance, computational overhead, and privacy protection. Specifically, we examine the effectiveness of homomorphic encryption, differential privacy, and a custom-developed fast and secure homomorphic encryption (FAS) technique. These methods were tested using different image datasets in a federated learning scenario simulated in a cloud environment.
Our key contributions are as follows:
Implementation of Security Techniques: We implemented secure aggregation methods, including homomorphic encryption, differential privacy, and FAS, within a federated learning model.
Development of FAS: We propose a novel FAS technique combining selective encryption, differential noise addition, and bitwise scrambling to achieve high security with minimal computational overhead.
Performance Evaluation: We provide a comparative analysis of encryption techniques, highlighting their overhead, scalability, and resilience against model inversion attacks, using metrics like MSSIM and VIFP.
Practical Insights: We assess the trade-offs between privacy, performance, and scalability, providing actionable insights for deploying secure federated learning systems in real-world applications.
The results demonstrate that FAS achieves a balance between robust security and computational efficiency, outperforming traditional techniques in scenarios where real-time performance and scalability are critical. The findings also highlight the suitability of FAS for resource-constrained environments like healthcare, where both privacy and efficiency are paramount.
The rest of this paper is organized as follows: The Datasets section describes the different image datasets used in our study. In Background on Deep Learning Models and Security Techniques, we discuss the deep learning models evaluated in our experiments, including ResNet-50 [2], DenseNet121 [3], EfficientNetB0 [4], MobileNet V2 [5], and elaborate on the security techniques applied.
The Experimental Setup for Federated Learning details our use of CloudLab [10] as the testbed and the Flower framework [11] for federated learning experiments. The Evaluation section presents encryption overhead and time costs across datasets and models, with a detailed time comparison of security techniques.
In Comparison of Security Techniques, we evaluate differential privacy [15], homomorphic encryption [16], and FAS, and compare these with similar works. The Practical Implications section discusses efficiency considerations and the balance between security and performance, followed by a Visual Comparison of FAS Approaches. Finally, we conclude with a summary of our findings and future research directions.
II Background and Related Work
Privacy-preserving techniques in federated learning have gained attention due to the need to protect sensitive data across distributed devices. Federated learning, introduced by McMahan et al. [23], enables decentralized model training but poses risks of data leakage through model updates.
Homomorphic encryption (HE) offers a robust solution, allowing computations on encrypted data while maintaining confidentiality. Gentry’s fully homomorphic encryption (FHE) [16] made arbitrary encrypted computations feasible, though its high computational overhead limits practical use [17, 28].
II-A Efficiency Enhancements in Homomorphic Encryption
Several advancements have sought to improve the efficiency of homomorphic encryption:
Fan and Vercauteren’s BGV Scheme [25]: This scheme optimizes the bootstrapping process, making HE more practical by efficiently handling both addition and multiplication on ciphertexts. This approach has been critical in making HE viable for more complex computations in federated learning environments [33]. Xu et al. [41] proposed HybridAlpha, an efficient privacy-preserving federated learning system that optimizes both privacy and performance. In comparison to our FAS approach, HybridAlpha leverages different privacy mechanisms such as differential privacy and secure aggregation to ensure efficient federated learning while maintaining data confidentiality.
Chillotti et al.’s TFHE Scheme [26]: TFHE introduced a faster bootstrapping process, making it highly efficient for secure Boolean computations, thus expanding the potential use of FHE in time-sensitive environments [34].
Ducas and Micciancio’s FHEW Scheme [27]: FHEW further improved the speed of homomorphic operations by optimizing the bootstrapping process, making FHE more feasible for larger-scale applications [27].
Cheon et al.’s CKKS Scheme [29]: This scheme supports approximate arithmetic operations on encrypted data, making it particularly efficient for machine learning and statistical analysis tasks where exact precision is not always required [36].
Brakerski, Gentry, and Vaikuntanathan’s BGV+ Scheme [28]: This work built on the BGV scheme, introducing further optimizations that reduce the computational overhead of FHE, particularly in noise management, which is crucial for maintaining the accuracy of federated learning models under encryption [28].
II-B Differential Privacy (DP) in Federated Learning
Differential privacy (DP) mitigates privacy risks by injecting controlled noise into model updates. Initially proposed by Dwork et al. [30], its integration into deep learning was further developed by Abadi et al. [31]. Subsequent work has refined DP methods to balance privacy and model accuracy [35].
II-C Selective Encryption in Federated Learning
Selective encryption reduces computational burden by encrypting only a portion of model parameters while maintaining security. Wu et al. [32] demonstrated its effectiveness in reducing overhead without compromising privacy. Similar approaches focus on optimizing encryption efficiency in federated learning [37, 38]. Zhang et al. [45] proposed BatchCrypt, an efficient homomorphic encryption framework that batches model updates to reduce the computational and communication overhead in federated learning. While BatchCrypt improves efficiency compared to traditional HE-based methods, it still incurs non-trivial overhead. Our approach further optimizes encryption by selectively encrypting a fraction of updates, achieving similar security guarantees with significantly lower computational costs. This paper extends these concepts by integrating selective encryption with differential noise and bitwise scrambling, reducing homomorphic encryption overhead while enhancing privacy protection.
II-D Bitwise Scrambling for Federated Learning Security
Bitwise scrambling further obfuscates partially encrypted data by rearranging bits using a cryptographic key. This technique complicates adversarial analysis and enhances security without significant computational cost.
Yang et al. [24] applied bitwise scrambling to secure data transmission, effectively preventing direct reconstruction of encrypted models. Halevi et al. [39] demonstrated its use with HE to resist chosen-ciphertext attacks while maintaining efficiency.
Combining selective encryption, differential privacy, and bitwise scrambling offers an effective privacy-preserving solution for federated learning, balancing security and computational efficiency for resource-constrained applications.
III Motivation and Threat Model
III-A Motivation
The adoption of federated learning (FL) in sensitive domains such as healthcare and finance necessitates robust privacy-preserving mechanisms. While FL mitigates data exposure by keeping raw data localized, model updates remain vulnerable to attacks like model inversion and gradient leakage, which can reveal sensitive information such as patient records or financial transactions.
Existing privacy-preserving techniques, such as fully homomorphic encryption (FHE) and differential privacy (DP), have limitations. FHE ensures confidentiality but incurs significant computational and communication overhead, making it impractical for resource-constrained settings. DP, though computationally efficient, can degrade model accuracy by introducing noise, which is problematic in high-precision applications.
To address these challenges, this work proposes the FAS approach, which balances security and efficiency by selectively encrypting critical model updates while applying differential noise and bitwise scrambling to the remaining data. This reduces computational overhead while maintaining strong privacy guarantees, making it particularly suitable for federated learning in time-sensitive applications like medical imaging and real-time analytics.
III-B Threat Model
Federated learning systems are susceptible to privacy threats due to decentralized training and the exchange of model updates. A semi-honest adversary may follow the protocol but attempt to infer sensitive information from intercepted or compromised data.
Adversaries can access model updates between clients and the central server, potentially compromising either to extract insights. Common attack vectors include gradient inversion to reconstruct training data, membership inference to verify specific data participation, and statistical analysis to infer private attributes. While SSL encryption secures transmissions, it does not prevent adversaries from analyzing transmitted updates.
To mitigate these risks, our approach integrates selective encryption, differential noise addition, and bitwise scrambling. Critical model parameters are encrypted, while unencrypted data is obfuscated with noise and scrambling to prevent reconstruction. Secure aggregation further ensures data protection throughout training, providing strong defenses against privacy attacks without incurring excessive computational overhead. This layered strategy effectively balances security and efficiency, making it well-suited for sensitive applications like healthcare.
IV Proposed Model
This research introduces a novel FAS (fast and secure federated learning) approach tailored for secure and efficient federated learning. The model aims to strike a balance between robust privacy preservation and computational efficiency, addressing the limitations of fully homomorphic encryption (high overhead) and differential privacy (accuracy degradation). Our approach integrates three key components: selective encryption, differential noise addition, and bitwise scrambling, providing a multi-layered defense against privacy attacks while maintaining high performance.
IV-A Selective Encryption
Rather than encrypting the entire model, FAS selectively encrypts a fixed percentage of critical model parameters using a homomorphic encryption scheme. This subset is chosen uniformly, ensuring that sensitive components are protected without the computational burden of encrypting the entire dataset. By encrypting only a fraction of the data, FAS significantly reduces the overhead associated with full encryption while maintaining strong security for the most sensitive parameters.
IV-B Differential Noise Addition
For the remaining unencrypted parameters, differential noise is added to obfuscate their true values. The noise is carefully calibrated using a Laplacian mechanism, ensuring that individual data points cannot be inferred while preserving the overall utility of the model. This approach leverages the principles of differential privacy to provide an additional layer of protection for unencrypted data.
IV-C Bitwise Scrambling
To further enhance privacy, unencrypted parameters are subjected to bitwise scrambling using a cryptographic key. This process disrupts patterns in the data, making it computationally infeasible for adversaries to reconstruct the original values without access to the scrambling key. By combining scrambling with noise addition, the model ensures that unencrypted data is protected against both statistical analysis and direct reconstruction attacks.
IV-D Federated Learning Integration
The FAS model integrates seamlessly into the federated learning workflow. Each client encrypts and obfuscates its local model updates using the proposed techniques before transmitting them to the central server. The server aggregates the updates without requiring decryption, preserving privacy throughout the computation. The aggregated model is then securely shared with clients for the next training iteration.
IV-E Security and Efficiency Trade-Off
The proposed model achieves a balance between security and efficiency:
  • Security: By encrypting critical parameters and obfuscating the rest, FAS provides strong resistance to model inversion and membership inference attacks, with security results comparable to full homomorphic encryption.
  • Efficiency: The selective encryption strategy reduces computational and communication overhead by up to 90%, making it suitable for resource-constrained federated learning environments.
This layered approach ensures that sensitive data remains protected while enabling efficient model training, making FAS an effective solution for privacy-preserving federated learning in applications like healthcare and finance.
IV-F Privacy-Preserving Techniques
IV-F1 Homomorphic Encryption
Homomorphic encryption (HE) is a cryptographic technique that allows computations to be performed directly on encrypted data without decrypting it. This ensures that sensitive information remains confidential throughout the computation process.
In the context of federated learning, HE enables secure aggregation of model updates from distributed clients. The mathematical foundation of HE ensures that:
$\text{Dec}(\text{Eval}(\text{Enc}(x),\text{Enc}(y)))=f(x,y),$
where Enc and Dec denote encryption and decryption, respectively, and Eval represents a computation on encrypted data. This property is critical for privacy-preserving applications, as it prevents adversaries from accessing plaintext data [16].
IV-F2 Differential Privacy
Differential privacy (DP) is a privacy-preserving framework that introduces controlled noise to statistical computations, ensuring that the presence or absence of a single data point does not significantly influence the output.
Formally, a mechanism $M$ satisfies $\epsilon$-differential privacy if for any two adjacent datasets $D_{1}$ and $D_{2}$ differing in a single entry, and for any subset of outputs $O$:
$\frac{\Pr[M(D_{1})\in O]}{\Pr[M(D_{2})\in O]}\leq e^{\epsilon},$
where $\epsilon$ is the privacy budget, controlling the trade-off between privacy and utility [30].
In federated learning, DP is used to add noise to model updates or gradients before aggregation, protecting individual contributions while preserving the overall utility of the model. This makes DP a valuable tool for maintaining privacy in sensitive domains such as healthcare and finance.
IV-G Process Flow of the Proposed Model
The proposed model operates in three main stages: client-side processing, server-side processing, and client post-processing. Each stage has a set of steps, illustrated in pseudocode and explained in detail below.
IV-H Federated Learning System Process Flow
Figure 2 illustrates the client-side process, while Figure 3 shows the server-side process in the federated learning system with encryption and scrambling techniques.
Figure 1: General depiction of the federated learning model showing the processes of encryption, scrambling, and aggregation.
Figure 1 provides a general depiction of our federated learning model, demonstrating the encryption, scrambling, and noise addition steps across client nodes, and the subsequent aggregation performed by the server.
Figure 2: Client-side process in federated learning system: encryption, scrambling, and noise addition.
Figure 3: Server-side process in federated learning system: aggregation and re-scrambling.
The process begins at the client side, where local data is used to train a model, producing a weight vector. This weight vector is split, encrypted, and further obfuscated through bitwise scrambling and the addition of noise.
Once encrypted, the weight vectors are formatted to hexadecimal and transmitted to the server. The server then receives encrypted data from all clients and performs federated averaging (FedAvg) to aggregate the model parameters. Importantly, the aggregation is done on both encrypted and non-encrypted vectors, ensuring privacy is maintained while also allowing for more efficient computation.
After aggregation, the server mixes the results with a scrambling key before sending the updated parameters back to the clients, where they are decrypted and applied to the local models for further training.
This secure and privacy-preserving process enables effective model training across decentralized data while preventing unauthorized access to sensitive information.
IV-H1 Client-Side Processing
In the client-side processing stage, model weights ($w$) are prepared for secure transmission to the server by encrypting selected weights, obfuscating the remaining weights, and combining them into a single data structure. Initially, the weights are split into two subsets based on a predefined encryption percentage ($enc\_pct$) (Lines 6–8). The selected portion of the weights is encrypted using a homomorphic encryption function ($enc\_func$), and these encrypted weights are stored in $enc\_w$ (Lines 9–11). For the remaining weights, they are first formatted to resemble encrypted data ($fmt\_weight$), noise is added using $noise\_func$, and the noisy weights are scrambled with a cryptographic key ($scr\_key$) via a scrambling function ($scr\_func$) (Lines 12–16). These scrambled weights are stored in $scr\_w$. Finally, the encrypted and scrambled weights are combined into a single structure ($enc\_scr\_w$) (Line 17) and transmitted to the server (Line 18).
Algorithm 1 Client-Side Enc, Scrambling, and Noise Addition
1:
2:$w$: Model weights
3:$enc\_pct$: Percentage of weights to encrypt
4:$enc\_func$: Encryption function
5:$scr\_key$: Scrambling key
6:$scr\_func$: Scrambling function
7:$noise\_func$: Noise addition function
8:
9:$enc\_scr\_w$: Processed (encrypted and scrambled) weights
10:procedure Client Enc-Scr-Noise
11:     Split $w$ into encrypted and non-encrypted parts
12:     $n\leftarrow$ length of $w$
13:     $num\_enc\leftarrow(enc\_pct/100)*n$
14:     $enc\_w\leftarrow[]$
15:     $scr\_w\leftarrow[]$
16:     for $i=0$ to $num\_enc-1$ do
17:         $enc\_weight\leftarrow enc\_func(w[i])$
18:         Append $enc\_weight$ to $enc\_w$
19:     end for
20:     for $i=num\_enc$ to $n-1$ do
21:         $fmt\_weight\leftarrow$ format_as_encrypted($w[i]$)
22:         $noisy\_weight\leftarrow noise\_func($fmt_weight))))
23:         $scr\_weight\leftarrow scr\_func($noisy_weight,scr_key))))
24:         Append $scr\_weight$ to $scr\_w$
25:     end for
26:     Combine $enc\_w$ and $scr\_w$ into $enc\_scr\_w$
27:     Send $enc\_scr\_w$ to the server
28:end procedure
IV-H2 Server-Side Processing
On the server side, the received combined weights ($enc\_scr\_w$) are processed to aggregate updates securely without decryption. The scrambled weights are first unscrambled using the same cryptographic key ($scr\_key$) applied during client-side processing (Lines 5–7). Once unscrambled, these weights are aggregated using a processing function ($srv\_proc\_func$) (Lines 8–9). Homomorphic operations are applied directly on the encrypted weights without decryption (Line 9). After processing, the unscrambled weights are reformatted and scrambled again using $scr\_func$ and $scr\_key$ to preserve security (Lines 10–13). These processed and re-scrambled weights are then combined with the processed encrypted weights and sent back to the client (Line 14).
Algorithm 2 Server-Side Processing without Decryption
1:
2:$enc\_scr\_w$: Received weights from client
3:$scr\_key$: Scrambling key for re-scrambling
4:$srv\_proc\_func$: Server processing function
5:
6:$sec\_w$: Processed weights sent back to client
7:procedure Server Proc w/o Decrypt
8:     Receive $enc\_scr\_w$ from the client
9:     $proc\_w\leftarrow[]$
10:     $unscr\_w\leftarrow[]$
11:     for $w$ in $scr\_w$ do
12:         $unscr\_weight\leftarrow unscramble\_func(w,scr\_key)$
13:         Append $unscr\_weight$ to $unscr\_w$
14:     end for
15:     Process $unscr\_w$ using $srv\_proc\_func$
16:     Process $enc\_w$ using homomorphic operations (no decryption needed)
17:     Format $unscr\_w$ back to scrambled format:
18:     for $w$ in $unscr\_w$ do
19:         $re\_scr\_w\leftarrow scr\_func(w,scr\_key)$
20:         Append $re\_scr\_w$ to $proc\_w$
21:     end for
22:     Combine $proc\_enc\_w$ with $re\_scr\_w$
23:     Send $sec\_w$ back to the client
24:end procedure
IV-H3 Client Post-Processing
In the final stage, the client receives the processed weights ($sec\_w$) from the server (Line 2) and restores them for updating the local model. The encrypted subset of weights is decrypted using the appropriate decryption function (Lines 3–4), while the scrambled weights are unscrambled with the cryptographic key ($scr\_key$) to restore their original form (Lines 5–7). These two subsets of weights are then combined to reconstruct the full model weights (Line 8), which are subsequently used to update the client’s local model for the next round of federated learning (Line 9).
Algorithm 3 Client Post-Processing with Noise Restoration
1:
2:$sec\_w$: Weights received from server
3:$scr\_key$: Scrambling key for restoration
4:
5:$final\_w$: Updated model weights for next round
6:procedure Client Post-Proc with Noise Rest
7:     Receive $sec\_w$ from the server
8:     $proc\_enc\_w\leftarrow sec\_w[0:num\_enc]$
9:     $proc\_scr\_w\leftarrow sec\_w[num\_enc:]$
10:     for $w$ in $proc\_scr\_w$ do
11:         $restored\_w\leftarrow reverse\_scr\_func(w,scr\_key)$
12:         Append $restored\_w$ to $rest\_w$
13:     end for
14:     Combine $proc\_enc\_w$ and $rest\_w$ to form $final\_w$
15:     Proceed to next round of federated learning with $final\_w$
16:end procedure
IV-I Advantages of the Proposed Model
  • Enhanced Security: The combination of encryption, noise, and scrambling protects both encrypted and unencrypted data.
  • Efficiency: Reduces computational overhead by selectively encrypting only a portion of the weights.
  • Scalability: Optimized for large-scale federated learning systems with multiple clients.
IV-J Privacy Analysis
In this section, we formally show that our federated learning mechanism, which combines selective homomorphic encryption, differential privacy, and bitwise scrambling, maintains a well-defined differential privacy guarantee.
IV-J1 Setup and Definitions
We consider a federated learning setting with $n$ clients, each holding private data. The global model is described by parameters indexed by $[N]$. Let $S\subseteq[N]$ be the subset of parameters to be protected by selective homomorphic encryption (HE), and let $[N]\setminus S$ be the remaining parameters to which we add differential privacy (DP) noise.
Differential Privacy.
A mechanism $M$ is $\epsilon$-DP if for any two adjacent datasets $D,D^{\prime}$ (differing in one record) and any event $O$:
$\frac{\Pr[M(D)\in O]}{\Pr[M(D^{\prime})\in O]}\leq e^{\epsilon}.$
[30]
Selective Homomorphic Encryption.
Parameters in $S$ are encrypted using a semantically secure HE scheme. Semantic security ensures no polynomial-time adversary can distinguish ciphertexts of different messages, implying no additional privacy cost in a DP sense (0-DP).[16]
Differential Privacy on the Remaining Parameters.
For each $i\in[N]\setminus S$, we add noise calibrated to $\epsilon_{i}$-DP. By the composition property of DP, releasing all these parameters together is $\left(\sum_{i\in[N]\setminus S}\epsilon_{i}\right)$-DP.[30]
Bitwise Scrambling.
We define a scrambling function $\mathcal{T}_{k}:\mathcal{X}\to\mathcal{X}$, where $\mathcal{X}$ is the space of possible model outputs. $\mathcal{T}_{k}$ is deterministic, keyed by $k$, and independent of $D$ except through $M(D)$. This scrambling is a form of post-processing.[43]
IV-J2 Privacy Guarantee
Theorem 1.
Consider a mechanism $\mathcal{M}$ that on input dataset $D$:
  1. 1.
    Encrypts $W_{i}$ for $i\in S$ using a semantically secure HE scheme [16, 44].
  2. 2.
    Adds noise to each $W_{j}$ for $j\in[N]\setminus S$ to achieve $\epsilon_{j}$-DP. Releasing all noisy parameters together is $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP [42].
  3. 3.
    Applies a scrambling function $\mathcal{T}_{k}$ to the entire output, resulting in $\mathcal{M}^{\prime}(D)=\mathcal{T}_{k}(\mathcal{M}(D))$ [43].
Then, $\mathcal{M}^{\prime}(D)$ is also $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP.
Proof.
Step 1 (DP on $[N]\setminus S$): For $[N]\setminus S$, each parameter $W_{j}+\text{noise}_{j}$ is $\epsilon_{j}$-DP. By composition, releasing all these parameters is $\sum_{j\in[N]\setminus S}\epsilon_{j}$-DP [42]. Formally, for any adjacent $D,D^{\prime}$ and event $O$:
$\frac{\Pr[\mathcal{M}_{[N]\setminus S}(D)\in O]}{\Pr[\mathcal{M}_{[N]\setminus S}(D^{\prime})\in O]}\leq e^{\sum_{j\in[N]\setminus S}\epsilon_{j}}.$
[42].
Step 2 (Selective Encryption on $S$): Parameters in $S$ are encrypted with a semantically secure HE scheme. Without the secret key, no information about the plaintext is revealed, contributing 0-DP cost [16, 44].
Step 3 (Scrambling as Post-Processing): The scrambling $\mathcal{T}_{k}$ is a deterministic, data-independent post-processing step. A fundamental property of DP is closure under post-processing: if $\mathcal{M}$ is $\epsilon$-DP, then $f(\mathcal{M})$ is also $\epsilon$-DP for any deterministic $f$ [43, 42]. Since $\mathcal{T}_{k}$ does not depend on $D$ beyond $\mathcal{M}(D)$ itself, applying $\mathcal{T}_{k}$ cannot increase or decrease the DP budget.
Combining these results:
$\frac{\Pr[\mathcal{M}^{\prime}(D)\in O]}{\Pr[\mathcal{M}^{\prime}(D^{\prime})\in O]}=\frac{\Pr[\mathcal{T}_{k}(\mathcal{M}(D))\in O]}{\Pr[\mathcal{T}_{k}(\mathcal{M}(D^{\prime}))\in O]}\leq e^{\sum_{j\in[N]\setminus S}\epsilon_{j}}.$
Thus, $\mathcal{M}^{\prime}(D)$ maintains the same DP guarantee as $\mathcal{M}(D)$. ∎
V Evaluation
This section evaluates our federated learning setup using various security techniques, including homomorphic encryption, differential privacy, and our proposed FAS. We assess their impact on test accuracy, communication cost, computational overhead, and privacy preservation.
Each dataset was split into 10 equal partitions, distributed across machines running Flower clients. Each machine trained its local model over 10 rounds, with 20 epochs per round. A validation set (10% of training data) ensured consistency, and final accuracy was evaluated on a separate test set. All models were trained using CPU resources to simulate realistic constraints.
V-A Datasets Used in Experiments
We selected publicly available datasets covering diverse imaging contexts with high usability ratings. Our criteria included broad disease coverage, diverse imaging techniques, and sufficient training and validation data.
V-A1 CIFAR-10 Dataset
A benchmark for image classification, CIFAR-10 consists of 60,000 32x32 images across 10 classes. It includes 50,000 training and 10,000 test images, commonly used for model performance evaluation [46].
V-A2 Chest X-ray (COVID-19, Pneumonia) Dataset
This dataset contains 6,339 grayscale X-ray images across three classes: benign, COVID-19, and pneumonia (2,313 images per class). These images aid in lung condition assessment [7].
Refer to caption Refer to caption Refer to caption
(a) Benign (b) COVID-19 (c) Pneumonia
Figure 4: Examples from the Chest X-ray (Covid, Pneumonia) dataset [20]
V-A3 CT Kidney Dataset
Comprising 12,446 CT scans categorized as benign (5,077), cyst (3,709), stone (1,377), and tumor (2,283), this dataset aids in kidney disease diagnosis [9].
Refer to caption Refer to caption
(a) Benign (b) Cyst
Refer to caption Refer to caption
(c) Stone (d) Tumor
Figure 5: Examples from the CT Kidney dataset [20]
V-A4 Diabetic Retinopathy Dataset
This dataset comprises 88,645 fundus images, categorized as benign (65,342) or diabetic retinopathy (23,303). Fundus imaging assists in detecting optic nerve abnormalities [6].
Refer to caption Refer to caption
(a) Benign (b) Diabetic retinopathy
Figure 6: Examples from the Diabetic Retinopathy Detection dataset [20]
V-A5 Lung Cancer Histopathological Images
This dataset contains 15,000 histopathological images, equally distributed across benign, squamous cell carcinoma (SCC), and adenocarcinoma (ACC) classes [8].
Refer to caption Refer to caption Refer to caption
(a) Benign (b) ACC (c) SCC
Figure 7: Examples from the Lung Cancer Histopathological Images dataset. [20]
Our evaluation primarily focused on the CT Kidney dataset, chosen for its stable, reliable results. The Chest X-ray and Diabetic Retinopathy datasets provided additional perspectives to assess model generalizability across different medical imaging contexts. The CT Kidney dataset’s robustness enabled prioritization of security technique evaluation over model-specific tuning.
VI Experimental Setup for Federated Learning
This section details the experimental setup for evaluating various cryptographic techniques in a federated learning context on medical image datasets.
VI-A CloudLab Environment
Experiments were conducted on CloudLab [10], a cloud computing testbed for systems research. Our setup utilized 11 physical machines in a shared-nothing cluster, connected via 10 Gbps Ethernet. Each machine was equipped with dual Intel E5-2683 v3 CPUs (14-core, 2.00 GHz), 256 GB RAM, and two 1 TB hard drives, providing substantial computational resources for federated learning tasks.
VI-B Federated Learning Framework: Flower
Flower [11], an open-source federated learning framework, was used for model training and evaluation, supporting TensorFlow [12], PyTorch [13], and MXNet [14]. One CloudLab machine served as the Flower server, while the remaining 10 functioned as Flower clients.
VI-C Federated Learning Setup
Each dataset was divided into 10 equal partitions, distributed across the Flower clients. The centralized federated learning setup comprised 10 clients and a server. Each model was trained over 10 rounds, with 20 epochs per round. A validation set (10% of training data) monitored performance, and final accuracy was evaluated on a test set. All training used CPUs to simulate resource-constrained environments.
VI-D Rationale for Centralized Federated Learning
Decentralized federated learning was not used due to a lack of mutual trust among clients. A centralized setup ensures data privacy, as clients communicate only with the server, avoiding direct data exchange and aligning with confidentiality requirements.
VI-E Deep Learning Models
We evaluated four deep learning models:
VI-E1 ResNet-50
ResNet-50 mitigates the vanishing gradient problem with skip connections, facilitating deep network training.
VI-E2 DenseNet121
DenseNet121 enhances feature reuse by connecting each layer to all previous layers, improving efficiency in resource-constrained settings.
VI-E3 EfficientNetB0
EfficientNetB0 optimally scales network width, depth, and resolution, balancing accuracy and efficiency.
VI-E4 MobileNet V2
MobileNet V2 employs lightweight convolutions to reduce computational costs while maintaining accuracy, making it suitable for mobile and embedded applications.
VI-F Encryption Configurations
We compared three encryption configurations:
  • Non-Enc: No encryption applied.
  • Full-Enc: All model weights encrypted using homomorphic encryption.
  • Partly-Enc: Selective encryption applied to a subset of data, with scrambling for security.
Full Overhead represents additional training time due to full encryption, while Partly Overhead captures the impact of partial encryption.
TABLE I: Training Time (Minutes) and Overheads Across Datasets
Model Dataset Full Enc (min) Partly Enc (min)
EffNetB0 CIFAR-10 111 21
CT Kidney 228 127
Lung 355 254
COVID 698 591
Diabetic Retinopathy 1298 1351
DenseNet121 CIFAR-10 219 35
CT Kidney 514 326
Lung 506 312
COVID 1542 1351
Diabetic Retinopathy 1542 1351
MobileNetV2 CIFAR-10 68 15
CT Kidney 163 100
Lung 467 231
COVID 610 555
Diabetic Retinopathy 610 555
ResNet-50 CIFAR-10 530 68
CT Kidney 861 399
Lung 834 372
COVID 1140 673
Diabetic Retinopathy 1140 673
Table I presents training times (minutes) across datasets. Fully encrypted models require significantly more training time, with ResNet-50 showing the highest overhead. In contrast, partly encrypted models reduce overhead substantially while maintaining security.
Figure 8: Encrypted file size per encryption level (10% - 100%)
Figure 8 shows communication costs across encryption levels, revealing higher costs for fully encrypted models, whereas partly encrypted models optimize performance.
Figure 9: Accuracy Vs Encryption Percentage Across Different Models
Figure 9 indicates that accuracy remains stable across encryption percentages, confirming the effectiveness of partly encryption in preserving model performance.
Figure 10: Training Time and Encrypted Size Comparison for Fully and Partly Encrypted Models (CT Kidney dataset)
Figure 10 highlights significant reductions in encryption overhead when using partly encryption, making FAS a viable option for resource-limited federated learning scenarios.
VI-G Effects on Security
Before conducting our main security experiments, we implemented several security tests to evaluate the robustness of our encryption methods. These tests included integrity checks, timing attack tests, noise and precision evaluations, and simulations of chosen-ciphertext and chosen-plaintext attacks. Additional differential attack tests and statistical distribution checks were performed to verify that encryption preserved data integrity and privacy. Our findings indicate that 10% partly encryption successfully passed the same security tests as 100% full encryption, demonstrating strong security with reduced computational overhead.
TABLE II: Full vs Partly Encryption: Security Test Results
Test Partly Enc. Full Enc.
Integrity Check False False
Chosen-Plaintext Attack True True
Chosen-Ciphertext Attack True True
Noise and Precision Test True True
Timing Attack Test True True
Differential Attack Test True True
Statistical Dist. Test True True
Homomorphism Test True True
Leakage Test True True
Table II compares partly encryption with full encryption across a series of security evaluations. The results confirm that partly encryption provides comparable protection, reinforcing its viability as a lightweight yet effective security mechanism.
VI-G1 Evaluating Privacy with MSSIM and VIFP
Model inversion attacks pose a threat to federated learning by reconstructing training data from model updates. To assess privacy resilience, we use Mean Structural Similarity Index (MSSIM) and Visual Information Fidelity in the Pixel domain (VIFP).
MSSIM quantifies structural similarity between original and reconstructed images, with lower scores indicating stronger privacy. VIFP measures visual quality based on perceptual models, highlighting the impact of noise and scrambling techniques. Both metrics are widely recognized for privacy evaluation in adversarial scenarios.
Our results show that selective encryption with scrambling and noise significantly degrades reconstruction quality. MSSIM stabilizes at 52% for 20% encryption, and VIFP follows a similar trend across multiple encryption levels. This confirms that our approach effectively mitigates inversion attacks while optimizing computational efficiency.
Further tests demonstrated that even at 10% encryption, reconstructed images showed significant distortion. The MSSIM score at 10% encryption was 58%, highlighting a substantial deviation from the original images. As encryption increased, the score stabilized at around 52%, reinforcing the effectiveness of partial encryption in maintaining privacy.
Figure 11: MSSIM and VIFP Scores across Encryption Percentages for Various Models
Figure 11 illustrates MSSIM and VIFP scores across encryption percentages for various models. Results were averaged across multiple runs to account for variations introduced by selective encryption, scrambling, and noise. The sharp decline in scores at 10% encryption, with stabilization around 20%, suggests that lower encryption levels can provide security equivalent to full encryption while significantly reducing computational costs. Notably, resistance to inversion attacks at 100% encryption is nearly identical to that at 10% and 20%, highlighting the robustness of selective encryption.
VI-H Cross-testing with Related Work
VI-H1 Comparison with MASKCRYPT
MASKCRYPT [19] applies a gradient-guided encryption mask to selectively encrypt critical model weights, reducing communication overhead by up to 4.15× compared to full model encryption. However, its gradient mask is not always optimized for selecting the most sensitive gradients, leading to lower initial security scores. MASKCRYPT requires multiple rounds to stabilize encryption, as reflected in VIFP scores, which only stabilize by round 5. This indicates an adaptation period where security is suboptimal.
In contrast, FAS stabilizes earlier by combining selective homomorphic encryption, differential noise, and bitwise scrambling, encrypting a fixed percentage of weights without gradient sensitivity analysis or pre-training. This approach reduces processing time by 90% compared to full homomorphic encryption while maintaining strong resistance to model inversion attacks. Unlike MASKCRYPT, which recalculates sensitivity masks each round, FAS eliminates extra computational costs and scales effectively across federated networks.
VI-H2 Comparison with FedML-HE
FedML-HE [18] identifies and encrypts critical weights based on gradient sensitivity through a pre-training phase. Our recreated FedML-HE model provides insight into its behavior, though results may differ from the original implementation. While FedML-HE offers strong privacy protection, its pre-training phase introduces significant time and resource costs. Each client generates an encryption mask independently, which, while avoiding direct mask sharing, can lead to inconsistencies in aggregated model accuracy.
Our FAS method eliminates the need for pre-training and mask aggregation by encrypting a fixed percentage of weights with differential noise and bitwise scrambling. At 10% encryption, FAS achieves a MSSIM score of 58%, comparable to FedML-HE’s 52%, but with significantly lower computational costs. The lack of pre-training and per-client mask generation allows FAS to maintain efficiency and scalability in federated learning environments.
Figure 12: Comparison of MSSIM and VIFP Scores: FAS vs. FedML-HE vs. MASKCRYPT
Figure 12 compares FAS, FedML-HE, and MASKCRYPT across MSSIM and VIFP scores. At 10% encryption, FedML-HE shows slightly higher stability (MSSIM 52%), while FAS achieves 58%, and MASKCRYPT falls in between at 55%. As encryption increases to 20%, all methods converge with nearly identical security performance. Despite slight advantages for FedML-HE in lower encryption levels, FAS provides comparable security at significantly lower computational costs. MASKCRYPT’s need for sensitivity mask recalculation each round makes it less efficient for real-time scenarios. These findings position FAS as an optimal choice for scalable, privacy-sensitive federated learning applications requiring low-latency performance.
TABLE III: Comparison of Encryption Techniques for Different Models
Method Model Non Enc Total (m) Overhead (m)
FAS EffNetB0 56 69 13
FEDML-HE EffNetB0 56 99 43
MASKCRYPT EffNetB0 56 89 33
FAS DenseNet121 152 176 24
FEDML-HE DenseNet121 152 210 58
MASKCRYPT DenseNet121 152 200 48
FAS MobileNetV2 46 52 6
FEDML-HE MobileNetV2 46 69 23
MASKCRYPT MobileNetV2 46 64 18
FAS ResNet-50 176 224 48
FEDML-HE ResNet-50 176 278 102
MASKCRYPT ResNet-50 176 246 70
Table III compares non-encrypted, partly encrypted, and overhead times for different models using FAS, FedML-HE, and MASKCRYPT on the Kidney dataset. The values are reconstructed based on methodologies described in the respective papers. FAS consistently has the lowest overhead across models, making it the most efficient. FedML-HE has the highest overhead due to pre-training, while MASKCRYPT offers a balance but initially has lower security, improving after multiple rounds. Overall, FAS is the best option for efficiency and scalability in federated learning.
VI-I Comparative Analysis of Encryption Techniques Across Datasets
This section compares the performance of FAS, FEDML-HE, and MASKCRYPT across five datasets: CIFAR-10, Diabetic Retinopathy, COVID, Lung, and Kidney. The analysis highlights FAS’s efficiency in minimizing overhead and encryption time compared to the other methods.
Across all datasets, FAS consistently achieves the lowest training times and computational overhead as we can see from Figures 13,  14,  17,  16,  15. Unlike FEDML-HE, which incurs significant pretraining costs for sensitivity mask creation, and MASKCRYPT, which generates masks at each round, FAS employs a random masking strategy with minimal computation. This advantage is particularly evident in lightweight models like MobileNetV2 on the COVID dataset and larger models like ResNet-50 on the Lung dataset. The efficiency of FAS is further demonstrated in the Diabetic Retinopathy and Kidney datasets, where it eliminates pretraining and per-round computations, making it the most practical and scalable approach. Overall, FAS outperforms the other techniques across all datasets, confirming its robustness and adaptability.
Figure 13: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (CIFAR-10).
Figure 14: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (Diabetic Retinopathy).
Figure 15: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (COVID).
Figure 16: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (Lung).
Figure 17: Comparison of Partly Encrypted and Fully Encrypted Metrics Across Models (Kidney).
Figures 13, 15, 14, 17 and 16 shows that across all datasets, FAS consistently outperforms FEDML-HE and MASKCRYPT in terms of total training time and overhead. The absence of sensitivity mask pretraining or extra per-round computations is the key factor contributing to its superior performance. These results establish FAS as the optimal choice for applications requiring fast and efficient encryption without compromising security. We can see that in smaller datasets which training time is not significant, difference between encryption technique becomes more obvious.
TABLE IV: Top 3 Cases Where FAS Outperforms Other Models
Dataset Model Compared To Enc. Improvement (%) Overhead Improvement (%)
COVID MobileNetV2 FEDML-HE 42.50 73.91
COVID EffNetB0 FEDML-HE 46.15 69.77
COVID EffNetB0 MASKCRYPT 36.36 60.61
Table IV presents the top three cases where the FAS encryption model demonstrates a higher efficiency compared to FEDML-HE and MASKCRYPT in terms of encryption time and overhead reduction.
COVID dataset with MobileNetV2: FAS achieves a 42.50% faster encryption time and a 73.91% lower overhead than FEDML-HE. COVID dataset with EffNetB0: FAS provides a 46.15% encryption improvement and a 69.77% overhead reduction compared to FEDML-HE. COVID dataset with EffNetB0 (vs. MASKCRYPT): FAS reduces encryption time by 36.36% and overhead by 60.61%. These results suggest that FAS is particularly effective in reducing computational overhead while maintaining faster encryption speeds, especially in the COVID dataset with lightweight models.
VI-J Handling Data Skew in Encryption-Based Techniques
In this section, we explore the impact of data skew on different encryption techniques, focusing on the robustness of the proposed FAS method compared to existing approaches such as MASKCRYPT and Fedml-HE.
VI-J1 Challenges with Mask-Based Techniques
Both MASKCRYPT and Fedml-HE rely on sensitive masks to select important gradients for encryption. These masks are designed to prioritize gradients that have a significant impact on the model’s accuracy. However, in scenarios with skewed data distributions, the importance of gradients becomes difficult to determine accurately. As a result, the encryption decisions made by these methods often degrade to levels resembling random encryption. This reduces their effectiveness, as observed in the MSSIM and VIFP score comparisons.
VI-J2 Robustness of the FAS Technique
In contrast, the proposed FAS technique, which operates as a random encryption method, is unaffected by data skew. Unlike mask-based techniques, FAS does not depend on the model’s accuracy or gradient importance for its encryption process. Instead, it employs a combination of encryption, scrambling, and noise addition. These components ensure that the method maintains consistent performance irrespective of the underlying data distribution. The robustness of this approach makes it particularly suitable for scenarios where data skew is prevalent, such as federated learning environments with heterogeneous clients.
VI-J3 Experimental Results and Comparisons
Dataset Method MSSIM (Skew/Normal) VIFP (Skew/Normal)
Kidney Fedml-HE 63 / 57 18 / 13
FAS 62 / 61 16.5 / 15
MASKCRYPT 63 / 60 18 / 15
Control 70 / 70 20 / 20
Lung Fedml-HE 65 / 59 19 / 14
FAS 62 / 61 16.5 / 15
MASKCRYPT 66 / 62 19 / 16
Control 70 / 70 20 / 20
COVID Fedml-HE 66 / 60 20 / 15
FAS 62 / 61 16.5 / 15
MASKCRYPT 67 / 63 20 / 17
Control 70 / 70 20 / 20
Diabetic Retinopathy Fedml-HE 64 / 58 18 / 14
FAS 62 / 61 16.5 / 15
MASKCRYPT 65 / 61 19 / 16
Control 70 / 70 20 / 20
CIFAR-10 Fedml-HE 62 / 58 17 / 13
FAS 61 / 59 16.5 / 15
MASKCRYPT 64 / 59 18 / 14
Control 70 / 70 20 / 20
TABLE V: MSSIM and VIFP Scores for Skew and Normal Data across Datasets
Table V demonstrates that the proposed FAS technique, leveraging encryption, scrambling, and noise, exhibits superior robustness to data skew across all evaluated datasets—Kidney, Lung, COVID, and Diabetic Retinopathy—when compared to existing gradient-based encryption methods. This characteristic ensures that FAS maintains its performance and security advantages even in diverse, real-world scenarios with varying data distributions. Experimental results using the EfficientNetB0 model across these datasets show that the accuracy of FAS is minimally affected by data skew, while other methods experience significant performance degradation under similar conditions. These findings highlight the adaptability and reliability of the FAS technique in privacy-preserving machine learning applications.
  • For MASKCRYPT and Fedml-HE, the MSSIM and VIFP scores under skewed data are significantl degraded, closely resembling random encryption.
  • In contrast, the FAS technique maintains reduced degregation across all conditions, as its random encryption approach is independent of the data distribution.
  • The control value has been kept same with skewed and normal to show the random selective encryption comparison.
Effect of Data Skew on Training Performance
Figure 18: Effect of Data Skew on CIFAR Dataset Across Different Techniques
Figure 18 illustrates the uniform reduction in accuracy across three techniques—FAS, MASKCRYPT, and Fedml-HE—when training on skewed data using the CIFAR dataset. While all techniques show similar accuracy drops under skewed conditions, the implications for their security mechanisms differ significantly.
  • MASKCRYPT and Fedml-HE: Both techniques rely on sensitivity masks to ensure data security. As training accuracy decreases under skewed conditions, the quality and reliability of these sensitivity masks are compromised, potentially leaving the system insecure until the mask stabilizes.
  • FAS: Unlike the other techniques, FAS does not depend on sensitivity masks. This independence ensures that the MS-SSIM and VIFP scores of FAS remain stable, even when training on skewed data. The inherent robustness of FAS allows it to maintain its security guarantees regardless of data distribution.
These results highlight a critical advantage of FAS: its ability to decouple data security from training accuracy. While longer training epochs can partially recover accuracy for MASKCRYPT and Fedml-HE, their reliance on sensitivity masks introduces a window of vulnerability during the calibration phase. In contrast, FAS maintains consistent security and performance, making it a more robust choice under challenging data conditions.
VII Conclusion
This paper evaluated privacy-preserving techniques in federated learning, focusing on different datasets and comparing differential privacy, homomorphic encryption, and our custom FAS approach.
Full homomorphic encryption offers the highest data protection but incurs significant computational costs, making it suitable only for scenarios prioritizing confidentiality over performance. Differential privacy provides lightweight privacy with minimal impact on computation, ideal for moderate security requirements. Our FAS method strikes a balance by achieving security comparable to full encryption while significantly reducing training time and overhead, making it efficient for large-scale federated learning and resource-constrained environments.
FAS’s layered approach—combining selective encryption, bitwise scrambling, and differential noise—demonstrates strong resilience against model inversion attacks without requiring pre-training or complex mask aggregation, outperforming FedML-HE and MASKCRYPT in scalability and efficiency. FAS offers slighly better security compared to models that require accurate sensitivity masks for data skews or operate under general low-accuracy conditions.
In summary, FAS offers an effective middle ground, balancing security and performance for real-time, privacy-sensitive applications like healthcare. Future research will refine these techniques and explore hybrid approaches across diverse datasets and federated environments to enhance scalability and applicability.
VIII Acknowledgments
The first author (A. K.) was supported by the Republic of Türkiye.